### v2

Abstract
Identifying low-dimensional structure in high-dimensional probability measures is an essential pre-processing step for efficient sampling. We introduce a method for identifying and approximating a target measure $\pi$ as a perturbation of a given reference measure $\mu$ along a few significant directions of $\mathbb{R}^{d}$. The reference measure can be a Gaussian or a nonlinear transformation of a Gaussian, as commonly arising in generative modeling. Our method extends prior work on minimizing majorizations of the Kullback–Leibler divergence to identify optimal approximations within this class of measures. Our main contribution unveils a connection between the dimensional logarithmic Sobolev inequality (LSI) and approximations with this ansatz. Specifically, when the target and reference are both Gaussian, we show that minimizing the dimensional LSI is equivalent to minimizing the KL divergence restricted to this ansatz. For general non-Gaussian measures, the dimensional LSI produces majorants that uniformly improve on previous majorants for gradient-based dimension reduction. We further demonstrate the applicability of this analysis to the squared Hellinger distance, where analogous reasoning shows that the dimensional Poincaré inequality offers improved bounds.
keywords:
Bayesian inference, dimensional logarithmic Sobolev inequality, dimensional Poincaré inequality, feature detection, gradient-based dimension reduction, reaction coordinates.
1 Introduction
Sampling from a given probability measure $\pi$ on $\mathbb{R}^{d}$ becomes increasingly challenging as the dimension $d$ increases [39, 38, 2, 35, 41, 40]. Efficient sampling algorithms therefore often rely on exploiting inherent low-dimensional structure present in $\pi$. Our motivation lies in identifying such a structure where $\pi$ is close, in a sense we make precise below, to another known probability measure $\mu$ on $\mathbb{R}^{d}$ which is itself easy to sample from. Typically, $\mu$ takes the form of a product measure, a Gaussian measure, or a nonlinear transformation of it, such as those used in GANs or other generative models [26]. Depending on the application, $\mu$ can be either given (e.g., via expert knowledge) or $\mu$ can constructed in some tractable class of measure (e.g., by minimizing some distance from $\pi$ to $\mu$).
In many realistic scenarios, $\pi$ is close to $\mu$ in that both measures essentially differ only along a few significant dimensions of $\mathbb{R}^{d}$. This suggests that $\pi$ can be effectively approximated by a probability measure $\widetilde{\pi}$ defined as
$\mathrm{d}\widetilde{\pi}(x)\propto\ell_{r}(U_{r}^{\top}x)\mathrm{d}\mu(x),$ (1.1)
where $\ell_{r}:\mathbb{R}^{r}\rightarrow\mathbb{R}_{+}$ is a low-dimensional function of $r\ll d$ variables and $U_{r}\in\mathbb{R}^{d\times r}$ a matrix, both of which need to be determined. We refer to ~$\widetilde{\ell}_{r}$ as the profile function and to $x\mapsto U_{r}^{\top}x$ as the linear features of $\widetilde{\pi}$. Such structured approximation — which we interchangeably describe as dimension reduction for the probability measure $\pi$ — is particularly relevant in at least the following contexts:
  • In Bayesian inverse problems [44, 31], $\pi$ can be the posterior measure defined as $\mathrm{d}\pi(x)\propto\ell(x)\mathrm{d}\mu(x)$, where $\mu$ denotes the prior measure of the model parameters $x\in\mathbb{R}^{d}$, and $\ell$ the likelihood function associated with observed data. The approximation (1.1) replaces the likelihood $\ell$ with a ridge function $x\mapsto\widetilde{\ell}_{r}(U_{r}^{\top}x)$, where the linear features $U_{r}$ thus capture low-dimensionality in the prior-to-posterior update of a Bayesian inverse problem (that is, the directions of the parameter $x$ for which data are most informative of).
  • In generative modeling, the probability density $\mu$ corresponding to empirical samples $\{X_{i}\}_{i=1}^{n_{\mathrm{train}}}$ in $\mathbb{R}^{d}$ is modeled as the push-forward of a lower-dimensional measure $\nu$ in $\mathbb{R}^{\kappa}$ with $\kappa\ll d$ latent coordinates. In other words, $\mu=T_{\#}\nu$, where the function $T:\mathbb{R}^{\kappa}\to\mathbb{R}^{d}$ is learned from samples [26]. It is of further interest to consider conditional sampling with a likelihood model specified in the nominal $d$ dimensional parameter space [37] — this is equivalent to sampling the Bayesian posterior $\mathrm{d}\pi(z)=(\ell\circ T)(z)\mathrm{d}\nu(z)$ in the latent space. The approximation (1.1) replaces the composed likelihood $\ell\circ T$ with a ridge function that captures low-dimensional prior-to-posterior updates, further complementing the dimensionality reduction brought on by the latent space representation.
  • In molecular dynamics, $\mathrm{d}\pi(x)\propto\exp(-\beta V(x))\mathrm{d}x$ can be a Boltzmann distribution associated with a molecular system with energy potential $x\mapsto V(x)$ with respect to the Lebesgue measure. Such distributions are typically multimodal, where each mode corresponds to a physical conformation. If a measure $\widetilde{\pi}$ as in (1.1) approximates well $\pi$ with $\mu$ being a unimodal measure (e.g. Gaussian), then features $U_{r}$ have given a physical interpretation as reaction coordinates which captures the multimodality of $\pi$. Finding good reaction coordinates is an active research field [36].
With some abuse of notation, we denote the measures in (1.1) by $\widetilde{\pi}(U_{r},\ell_{r})$. A natural idea for determining optimal approximations within this class is to consider minimizing an Amari $\alpha$-divergence [3]
$\min_{\begin{subarray}{c}U_{r}\in\mathbb{R}^{d\times r}\\ U_{r}^{\top}U_{r}=I_{r}\end{subarray}}\,\min_{\ell_{r}:\mathbb{R}^{r}\to\mathbb{R}_{+}}D_{\alpha}(\pi\,||\,\widetilde{\pi}(U_{r},\ell_{r})),$ (1.2)
where choosing $\alpha\in\mathbb{R}$ interpolates the Kullback-Leibler divergence ($\alpha=1$), the squared Hellinger distance ($\alpha=1/2$), the $\chi^{2}$ divergence ($\alpha=2$), and their reverse formulations. For any fixed $U_{r}$, the optimal profile function $\ell_{r}^{*}=\ell_{r}^{*}(U_{r})$ depends on $\alpha$ and $U_{r}$, but is known in closed-form, see e.g. [34, Thm 1]. Problem (1.2) thus reduces to optimizing $D_{\alpha}(\pi\,||\,\widetilde{\pi}(U_{r},\ell_{r}^{*}(U_{r})))$ for the features $U_{r}$. Nevertheless, this optimization requires access to (gradients of the) normalizing constants of $\widetilde{\pi}(U_{r},\ell_{r}^{*}(U_{r}))$ and $\pi$, which becomes computationally intractable for higher dimensional settings.
To circumvent this problem, Zahm et al. [48] proposed to majorize (1.2) with bounds of the form
$\min_{\ell_{r}:\mathbb{R}^{r}\to\mathbb{R}_{+}}D_{\alpha}(\pi\,||\,\widetilde{\pi}(U_{r},\ell_{r}))\leq\mathcal{J}_{\alpha}(U_{r}).$ (1.3)
The function $U_{r}\mapsto\mathcal{J}_{\alpha}(U_{r})$, which implicitly depends on $\pi$, should be easier to compute numerically and to minimize compared to the left-hand side of (1.3). Although minimizing this surrogate loss yields features $U_{r}^{*}$ which may be sub-optimal compared to the global minimizer of the $\alpha$-divergence, this procedure furnishes $\widetilde{\pi}(U_{r}^{*},\ell_{r}^{*}(U_{r}^{*}))$ with computationally certifiable approximation error. Accordingly, one can then select $r$ to achieve any desired error tolerance $D_{\alpha}(\pi\,||\,\widetilde{\pi}(U_{r}^{*},\ell_{r}^{*}(U_{r}^{*})))\leq\epsilon$ by ensuring $\mathcal{J}_{\alpha}(U_{r}^{*})\leq\epsilon$. In many instances it suffices to choose $r\ll d$ to achieve  $\epsilon\ll 1$.
The construction of these majorizations hinges on functional inequalities satisfied by the reference measure $\mu$, and only requires the Radon-Nikodym derivative $\mathrm{d}\pi/\mathrm{d}\mu$ to be a smooth function. The seminal contribution of Zahm et al. [48] related dimension reduction with the Kullback-Leibler (KL) divergence to the logarithmic Sobolev inequality (LSI, [28]), while afterwards [20, 17, 25] related the squared Hellinger distance to the Poincaré inequality. Both constructions are generalized in [34], which related each Amari $\alpha$-divergence to a corresponding $\phi$-Sobolev inequality, see [18, 6]. Remarkably, in all these works the minimizers $U_{r}^{*}$ of $U_{r}\mapsto\mathcal{J}_{\alpha}(U_{r})$ are identical, and are given by the matrix containing the dominant eigenvectors of
$H(\pi||\mu)=\mathbb{E}_{\pi}\left[\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}\mu}\right)\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}\mu}\right)^{\top}\right],$ (1.4)
which is the Fisher information matrix of $\pi$ relative to $\mu$. While these functional inequalities seemingly suggest that the relative Fisher information features are quasi-optimal for all $\alpha$, it remains unclear whether better majorants can be obtained for specific $\alpha$ using more specialized functional inequalities.
Our main contribution is to show that the dimensional logarithmic Sobolev inequality [10, 22] permits us to obtain better majorants for (and only for) the KL divergence corresponding to $\alpha=1$. Assuming $\mu$ to be the standard Gaussian measure in $\mathbb{R}^{d}$, this functional inequality improves upon the LSI by explicitly accounting for the dimension $d$, and formally recovers the latter in the limit $d\to\infty$. In our Theorem 4, we show how to exploit the dimensional LSI in order to obtain both a majorant $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ and a minorant $\mathcal{J}_{\mathrm{KL}}^{\uparrow}(\cdot)$ of the KL divergence so that
$\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})\leq\min_{\ell_{r}:\mathbb{R}^{r}\to\mathbb{R}_{+}}D_{\mathrm{KL}}(\pi\,||\,\widetilde{\pi}(U_{r},\ell_{r}))\leq\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r}).$ (1.5)
(We suppress their implicit dependence on $\pi$ for notational levity.) While the minorant $\mathcal{J}_{\mathrm{KL}}^{\uparrow}(\cdot)$ is not used to construct $U_{r}$, it provides a useful certificate of the sharpness of inequality (1.5) at minimizers of the majorant. Notably, the majorant $\mathcal{J}^{\downarrow}_{\mathrm{KL}}(\cdot)$ satisfies $\mathcal{J}^{\downarrow}_{\mathrm{KL}}(U_{r})\leq\mathcal{J}_{1}(U_{r})$ for all $U_{r}$ and thus uniformly improves upon the bound (1.3) obtained in [48]. The critical points of $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ and $\mathcal{J}_{1}(\cdot)$, however, do not necessarily coincide.
Heuristically, the motivation for pursuing this new choice of functional inequality stems from the observation that the LSI is satisfied by ‘infinite dimensional’ distributions (this is implied by standard tensorization arguments, see e.g. [8]). Since our intended application is dimension reduction, intuitively the dimensional LSI should be the more appropriate tool. In fact, we discover that the dimensional LSI instrinically captures dimension reduction for probability measures: our Proposition 7 shows that the dimensional majorant $\mathcal{J}^{\downarrow}_{\mathrm{KL}}(U_{r})$ is equivalent to the KL divergence (1.5) whenever $\pi$ is a Gaussian measure. When $\pi$ corresponds to a non-Gaussian measure, we show that the majorant in (1.5) is strictly tighter than the majorant obtained from the LSI, with the difference being exponentially better in certain instances.
Evaluating and optimizing the majorant $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ requires computing
$H(\pi)$ $=\mathbb{E}_{\pi}\left[\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)^{\top}\right]\quad\text{and}\quad M(\pi)=\mathbb{E}_{\pi}[XX^{\top}],$ (1.6)
where $H(\pi)$ is the Fisher information matrix of $\pi$, relative to the Lebesgue measure $\mathrm{d}x$ of $\mathbb{R}^{d}$, and $M(\pi)$ is the second moment matrix of $\pi$. Evaluating the minorant $\mathcal{J}_{\mathrm{KL}}^{\uparrow}(\cdot)$ requires in addition the computation of $m(\pi)=\mathbb{E}_{\pi}[X]$, the mean of the measure $\pi$. However, unlike for the LSI majorant $\mathcal{J}_{1}(\cdot)$, there is no closed-form expression for the global minimizer of $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$. Nevertheless, we observe numerically that first order gradient descent methods converge to the same minimizer irrespective of the initialization. We therefore conjecture — but cannot prove — that $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ is benignly non-convex [42] on the Grassmann manifold $\{U_{r}\in\mathbb{R}^{d\times r}:U_{r}^{\top}U_{r}=I_{r}\}$, and we leave this as a problem of independent interest left for future work.
We also consider adapting the reference measure $\mu$ to be the best Gaussian approximation to $\pi$, with mean $m(\pi)$ and covariance $C(\pi)$, given by $\mu=\mathcal{N}(m(\pi),\,C(\pi))$. In Theorem 8 we show that this choice, combined with the dimensional LSI, leads to a majorization whose global minimizer is known in closed form. Intriguingly, this bound suggests that relative deviations between $\pi$ and its best Gaussian approximation are best captured by the relative changes between the Fisher information matrix $H(\pi)$ and the precision matrix $C(\pi)^{-1}$, in that
$\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{r}^{\downarrow}))\leq\frac{1}{2}\sum_{k=r+1}^{d}\ln(\lambda_{k}(H(\pi),C(\pi)^{-1})),$ (1.7)
where $\lambda_{k}(A,B)$ is the $k$-th largest eigenvalue of the generalized eigenvalue problem $Av=\lambda Bv$, and the columns of $U_{r}^{\downarrow}$ contain the first $r$ generalized eigenvectors.
When considering Bayesian inverse problems specifically, [21] shows that the LSI can be used to determine features $V_{r}^{*}$ along which the posterior $\pi_{X\mid Y}$ differs most from the prior $\mu_{X}$ for the average realization of data $Y$. In Theorem 9 we show that the dimensional LSI further improves on these bounds. This recovers the same features in [21], with $V_{r}^{*}$ corresponding to the leading eigenvectors of $H_{\mathrm{df}}(\pi_{X\mid Y}):=\mathbb{E}_{X,Y}[\nabla_{x}\ln\pi_{x\mid y}\nabla_{x}\ln\pi_{x\mid y}^{\top}]$, but offers the significantly improved certificate
$\mathbb{E}_{Y}\left[\operatorname{D_{\mathrm{KL}}}(\pi_{X\mid Y}\,||\,\widetilde{\pi}_{X\mid Y}(V_{r}^{*})\right]\leq\frac{1}{2}\sum_{k=r+1}^{d}\ln(1+\lambda_{k}(H_{\mathrm{df}}(\pi_{X\mid Y})),$ (1.8)
where $\lambda_{k}(\cdot)$ denotes the $k$-th largest eigenvalue of $H_{\mathrm{df}}(\pi_{X\mid Y})$. Practically, (1.8) implies that significantly fewer features are necessary to attain a certifable error level than previously believed. The form of the certificate (1.8) also highlights a connection to the expected information gain (EIG) appearing in optimal experimental design [1, 29]. The application of functional inequalities to optimal experimental design is also a topic of recent interest [33, 16].
Lastly, we also explore structured approximations in (1.2) with the squared Hellinger distance corresponding to $\alpha=1/2$. Motivated by the same heuristic considerations behind exploring the dimensional LSI, we show in Theorem 13 that the dimensional Poincaré inequality proposed in [7] permits us to obtain improved majorizations over the bounds obtained from the standard Poincaré inequality. However, the form of this improved majorization is implicit, which is difficult to optimize directly. Instead, we consider its use in obtaining improved certificates for the optimal features $U_{r}^{*}$. Intriguingly, leveraging this improvement requires the use of lower bounds on the approximation error, which presents significant computational challenges. We leave addressing these issues, as well as potential applications of the dimensional Poincaré to more general ridge-based regression (see e.g. [12, 47]), to future work.
2 Dimension Reduction with the KL Divergence
2.1 Certifiable bounds using the logarithmic Sobolev inequality
We recall some properties of the (KL divergence) optimal profile function from [48, §2.1]. For any matrix $U_{r}\in\mathbb{R}^{d\times r}$ with $r\leq d$, and with $\widetilde{\pi}$ as in (1.1), the minimizer of $\ell_{r}\mapsto\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi})=\int\log(\frac{\mathrm{d}\pi}{\mathrm{d}\widetilde{\pi}})\mathrm{d}\pi$ is the conditional expectation
$\ell^{\mathrm{KL}}_{r}(\theta_{r})=\mathbb{E}_{X\sim\mu}\left[\left.\frac{\mathrm{d}\pi}{\mathrm{d}\mu}(X)\,\right|\,U_{r}^{\top}X=\theta_{r}\right].$ (2.1)
We denote by $\widetilde{\pi}^{\mathrm{KL}}(U_{r})=\widetilde{\pi}^{\mathrm{KL}}(\cdot\mid U_{r})$ the resulting probability measure given by
$\mathrm{d}\widetilde{\pi}^{\mathrm{KL}}(x\mid U_{r})\propto\ell^{\mathrm{KL}}_{r}(U_{r}^{\top}x)\mathrm{d}\mu(x),$ (2.2)
Let us emphasize that, by construction, the law of $U_{r}^{\top}\widetilde{X}$ and of $U_{r}^{\top}X$ are the same for any $U_{r}$, where $\widetilde{X}\sim\widetilde{\pi}^{\mathrm{KL}}_{r}(U_{r})$ and $X\sim\pi$. In other words, the KL optimal profile (2.1) ensures that the marginal laws of $\widetilde{\pi}^{\mathrm{KL}}_{r}(U_{r})$ and $\pi$ are the same along the linear feature $U_{r}$. Using the above notations, problem (1.2) becomes
$\min_{\begin{subarray}{c}U_{r}\in\mathbb{R}^{d\times r}\\ U_{r}^{\top}U_{r}=I_{r}\end{subarray}}\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r})).$ (2.3)
In the following we assume that $\mu=\mathcal{N}(0,I_{d})$ is the standard normal measure. In order to derive a majorant for $\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}))$, and we employ logarithmic Sobolev inequality (LSI).
Proposition 1 (Gaussian LSI).
The entropy $\textsc{Ent}_{\mu}(f)=\int f\ln f\mathrm{d}\mu-\int f\mathrm{d}\mu\ln\int f\mathrm{d}\mu$ of any smooth positive function $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$ with respect to the standard Gaussian $\mu=\mathcal{N}(0,I_{d})$ satisfies
$\frac{1}{2\int f\mathrm{d}\mu}\,\left\|\int\nabla f\mathrm{d}\mu\right\|_{2}^{2}\leq\textsc{Ent}_{\mu}(f)\leq\frac{1}{2}\int\|\nabla\ln f\|_{2}^{2}f\mathrm{d}\mu.$ (2.4)
We refer to the upper bound as the logarithmic Sobolev inequality (LSI), and to the lower bound as the reverse logarithmic Sobolev inequality (reverse LSI).
The logarithmic Sobolev inequality first appeared in the seminal paper of Gross [28] — we are unaware of the original reference for the reverse inequality. Both bounds are formulated for general Markov diffusion operators in [8, Ch. 5], which we translate in Appendix C.1 to obtain the above proposition. Note the choice of $f=\mathrm{d}\pi/\mathrm{d}\mu$ in (2.4) permits us to bound the divergence $\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\mu)$ following
$\frac{1}{2}\left\|\mathbb{E}_{\pi}\left[\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}\mu}\right)\right]\right\|_{2}^{2}\leq\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\mu)\leq\frac{1}{2}\mathbb{E}_{\pi}\left[\left\|\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}\mu}\right)\right\|_{2}^{2}\right],$
The next theorem further motivates the connection between dimension reduction for the KL divergence and the logarithmic Sobolev inequality, and we leave the proof to Appendix D.1. While this strategy was originally proposed in [48, Corr. 2.10], the formulation of the lower bound is a new contribution.
Theorem 2.
Consider the probability measure $\mathrm{d}\pi(x)\propto\ell(x)\mathrm{d}\mu(x)$ for some smooth $\ell:\mathbb{R}^{d}\to\mathbb{R}_{+}$ and standard Gaussian $\mu=\mathcal{N}(0,I_{d})$. Then, for any matrix $U_{r}\in\mathbb{R}^{d\times r}$ with $r\leq d$ orthonormal columns, the measure $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$, as in (2.2), satisfies
$\frac{1}{2}\|(I_{d}-U_{r}U_{r}^{\top})m(\pi)\|_{2}^{2}\leq\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}))\leq\frac{1}{2}\operatorname{tr\,}\left((I_{d}-U_{r}U_{r}^{\top})H(\pi||\mu)\right),$ (2.5)
where $m(\pi)=\mathbb{E}_{\pi}[X]$ is the mean of $\pi$ and $H(\pi||\mu)=\mathbb{E}_{X\sim\pi}[\nabla\ln\ell(X)\nabla\ln\ell(X)^{\top}]$ is the Fisher information matrix of $\pi$ relative to $\mu$.
Minimizing the right-hand side of (2.5) corresponds corresponds to maximimizing $U_{r}\mapsto\operatorname{tr\,}(U_{r}^{\top}H(\pi||\mu)U_{r})$ and yields the matrix $U_{r}^{*}$ whose columns contain the $r$ leading eigenvectors of $H(\pi||\mu)$, see [48]. This is computationally tractable to obtain, requiring only (an empirical estimate of) the relative Fisher information matrix. The approximation of $\pi$ by $\widetilde{\pi}^{\mathrm{KL}}(U_{r}^{*})$ is therefore furnished with the error certificates
$\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}^{*}))\leq\frac{1}{2}\sum_{k=r+1}^{d}\lambda_{k}(H(\pi||\mu)),$ (2.6)
where $\lambda_{k}(H)$ denotes the $k$-th largest eigenvalue of $H(\pi||\mu)$. While $U_{r}^{*}$ may not be the globally optimal minimum of (2.3), this result certifies that its approximation error is no larger than $\frac{1}{2}\sum_{k=r+1}^{d}\lambda_{k}(H(\pi||\mu))$, which can be made arbitrarily small as $r\to d$. When the spectrum of $H$ decays rapidly, it suffices to choose rank $r\ll d$ to obtain an accurate approximation. The lower bound (2.5) provides a guarantee of the smallest achievable approximation error, which may also inform the practitioner’s choice of $r$.
2.2 Improved bounds using dimensional inequalities
We now introduce a functional inequality which is uniformly sharper than the LSI for Gaussian measures. We defer its proof to Appendix C.2. In the following, we employ the notation $v^{\otimes 2}=vv^{T}$.
Proposition 3 (Dimensional Gaussian LSI).
For standard Gaussian $\mu=\mathcal{N}(0,I_{d})$ on $\mathbb{R}^{d}$, and for any smooth function $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\int f\mathrm{d}\mu=1$, we have the dimensional logarithmic Sobolev and the reverse dimensional logarithmic Sobolev inequalities
$\textsc{Ent}_{\mu}\left(f\right)$ $\leq{\frac{1}{2}\int\|x\|_{2}^{2}f\mathrm{d}\mu-\frac{d}{2}+\frac{1}{2}\ln\det\left(\int\left(\nabla\ln f-x\right)^{\otimes 2}f\mathrm{d}\mu\right),}$ (2.7)
$\textsc{Ent}_{\mu}(f)$ $\geq\frac{1}{2}\int\|x\|_{2}^{2}f\mathrm{d}\mu-\frac{d}{2}-\frac{1}{2}\ln\det\left(\int x^{\otimes 2}\,f\mathrm{d}\mu-\left(\int xf\mathrm{d}\mu\right)^{\otimes 2}\right).$ (2.8)
The Gaussian dimensional LSI was first proposed by Bakry and Ledoux [10, Prop. 2], but the formulations in Proposition 3 are attributed to Dembo [22] [^1] We were made aware of this reference from [24], but could not obtain a copy ourselves., who strengthened their proof technique in a natural manner (see Appendix C.2). By concavity $\ln\det(S)\leq\operatorname{tr\,}S-d$, which shows that (2.7) and (2.8) are uniformly tighter than the bounds in Proposition 1. Due to the explicit dependence on dimension, however, these inequalities do not tensorize. Note that the constants in Proposition 3 are sharp for functions given by (translations of) Gaussian kernels [10, §2], just as with the LSI.
Remark 1.
We are only aware of the following classes of measures satisfying the dimensional LSI in the Euclidean geometry: Gaussian measures (with possibly singular covariances), push-forwards of Gaussian measures by Lipschitz functions, and perturbations of Gaussians measures in the sense of Holley-Stroock [30]. Formulations of the dimensional logarithmic Sobolev inequalities exist for measures on other manifolds, but we leave the interested reader to e.g., [8, Thm 6.7.3] for further details.
Improved bounds for the KL divergence between $\pi$ and its approximation $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ are obtained using the dimensional LSI; we leave the proof to Appendix D.2.
Theorem 4.
Consider the standard Gaussian measure $\mu=\mathcal{N}(0,I_{d})$ on $\mathbb{R}^{d}$. Then, for any matrix $U_{r}\in\mathbb{R}^{d\times r}$ with $r\leq d$ orthonormal columns, the measure $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$, as in (2.2), satisfies
$\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})\leq D_{\mathrm{KL}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}))\leq\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r}),$ (2.9)
where
$\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})=\frac{\operatorname{tr\,}(M(\pi))-d+\ln\det\left(H(\pi)\right)}{2}-\frac{\operatorname{tr\,}(U_{r}^{\top}M(\pi)U_{r})-r+\ln\det(U_{r}^{\top}H(\pi)^{-1}U_{r})}{2},$ (2.10)
$\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})=\frac{\operatorname{tr\,}(M(\pi))-d-\ln\det(C(\pi))}{2}-\frac{\operatorname{tr\,}(U_{r}^{\top}M(\pi)U_{r})-r-\ln\det(U_{r}^{\top}C(\pi)^{-1}U_{r})}{2}.$ (2.11)
Here $H(\pi)=\mathbb{E}_{\pi}[\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}x})\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}x})^{\top}]$ denotes the Fisher information matrix of $\pi$ and $M(\pi)=\mathbb{E}_{\pi}[XX^{\top}]$ is the second moment matrix of $\pi$, and $C(\pi)=M(\pi)-\mathbb{E}_{\pi}[X]\mathbb{E}_{\pi}[X]^{\top}$ is the covariance matrix of $\pi$.
Remark 2.
Let $U_{\perp}$ denote any orthogonal completion to $U_{r}$ such that $[U_{r},U_{\perp}]\in\mathbb{R}^{d\times d}$ is unitary. Using standard matrix determinant identities (see Appendix B) we have the equivalent representations
$\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})$ $=\frac{\operatorname{tr\,}(U_{\perp}^{\top}M(\pi)U_{\perp})-(d-r)+\ln\det(U_{\perp}^{\top}H(\pi)U_{\perp})}{2}$ (2.12)
$\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})$ $=\frac{\operatorname{tr\,}(U_{\perp}^{\top}M(\pi)U_{\perp})-(d-r)-\ln\det(U_{\perp}^{\top}C(\pi)U_{\perp})}{2},$ (2.13)
which do not require computing matrix inverses to evaluate. Interestingly, the above expressions resemble, but fundamentally differ from, the log-determinant divergence (or Burg divergence) $D_{\ell,d}(X,Y)=\operatorname{tr\,}(XY^{-1})-d-\ln\det(XY^{-1})$ between two symmetric positive definite matrices $X,Y\in\mathbb{R}^{d\times d}$, see [32].
Just as the dimensional LSI improves on the LSI, the inequalities in Theorem 4 improve on those from Theorem 2, meaning that
$\frac{1}{2}\|(I_{d}-U_{r}U_{r}^{\top})m(\pi)\|_{2}^{2}\leq\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})\qquad\text{and}\qquad\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})\leq\frac{1}{2}\operatorname{tr\,}\left((I_{d}-U_{r}U_{r}^{\top})H(\pi||\mu)\right).$
To demonstrate this, it suffises to apply $\ln\det(S)\leq\operatorname{tr\,}(S)-d$ on (2.12) and (2.13) and observe that the relative Fisher information matrix $H(\pi||\mu)$ with $\mu=\mathcal{N}(0,I_{d})$ decomposes as
$H(\pi||\mu)$ =(1.4)${=}\mathbb{E}_{\pi}\left[\left(\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)+x\right)\left(\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)+x\right)^{\top}\right]$
=(1.6)${=}H(\pi)+\left(\int\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)x^{\top}+x\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\right)^{\top}\mathrm{d}\pi\right)+M(\pi)$ )
$=H(\pi)-2I_{d}+M(\pi),$ (2.14)
where the last step is obtained by an integration by parts. Theorem 4 sharpens the error certificates from Theorem 2 for all features by an exponential factor (the “logdet” term). Minimizing $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ thus results in features with tighter certifiable approximation guarantees compared to minimizing the majorant in (2.5), as shown in Appendix A.
We show in the following proposition that the majorization (2.10) monotonically decreases over sequences of nested features $\textrm{span}(U_{r})\subseteq\textrm{span}(U_{r+1})$. This implies $\min_{U_{r+1}}\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r+1})\leq\min_{U_{r}}\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})$, and thus one always obtains features with improved error certificates when increasing the feature rank $r$, as intuitively desired. We defer the proof to Appendix D.3. However, unlike the minimizers $\{U_{1}^{*},U_{2}^{*},\ldots\}$ of the LSI bound (2.5) which satisfy $\textrm{span}(U_{r}^{*})\subseteq\textrm{span}(U_{r+1}^{*})$, the minimizers of $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ are not necessarily nested.
Proposition 5.
Let $\{u_{k}\}_{k=1}^{d}$ denote any orthonormal basis of $\mathbb{R}^{d}$ and consider the sequence $\{U_{r}=(u_{1},\ldots,u_{r})\in\mathbb{R}^{d\times r}\}_{1\leq r\leq d}$. Then, we have $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r+1})\leq\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})$ for all $r<d$. As a consequence, for all $r<d$ we have
$\min_{\begin{subarray}{c}U_{r+1}\in\mathbb{R}^{d\times(r+1)}\\ U_{r+1}^{\top}U_{r+1}=I_{r+1}\end{subarray}}\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r+1})\leq\min_{\begin{subarray}{c}U_{r}\in\mathbb{R}^{d\times r}\\ U_{r}^{\top}U_{r}=I_{r}\end{subarray}}\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r}).$ (2.15)
There are additional numerical considerations to minimizing $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$, however. When viewed as a function over Euclidean vectors in $\mathbb{R}^{d\times r}$, this objective is non-convex due to the orthogonality constraint $U_{r}^{\top}U_{r}=I_{r}$. Alternatively, when considered as optimization over the Grassmann manifold $\mathrm{Gr}(d,r)$ — the manifold of $r$ dimensional subspaces equipped with the canonical Euclidean metric [23] — it is also impossible for this function to be globally geodesically convex, as such smooth functions cannot exist on compact manifolds [14, Corr 11.10]. Both observations suggest that (Riemannian) gradient descent may stagnate at local minima. Despite this, our numerical experiments in Pymanopt [46] — with structured $M(\pi)$ and $H(\pi)$ arising from statistical problems, or more generally with random $M(\pi)\succ 0$ and $H(\pi)\succ 0$ — suggests that Riemannian gradient descent [23] converges to the same solution for all initial conditions. This leads us to the following.
Conjecture 6.
The function $\bar{\mathcal{J}}_{\mathrm{KL}}^{\downarrow}:\mathrm{Gr}(d,r)\rightarrow\mathbb{R}$ defined by $\bar{\mathcal{J}}_{\mathrm{KL}}^{\downarrow}(\mathrm{range}\{U_{r}\})=\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})$ as in (2.10) for all orthonormal matrix $U_{r}\in\mathbb{R}^{d\times r}$, is benignly non-convex on the Grassmann manifold $\mathrm{Gr}(d,r)$.
Benign non-convexity refers to non-convex functions for which every local minima is a global minima; see e.g. [42, 45] for a curated list of examples. Geometrically, all saddle points (if any) must therefore have at least one escape direction, providing a mechanism for first-order descent methods to reach the global minimizer almost surely. While there has been significant interest in the optimization community on discovering such functions, to our knowledge the proof techniques for demonstrating benign non-convexity on the Grassmannian require knowing the global minima in closed form, see e.g. [4, 5]. We are unable to prove such a characterization, and we can only derive necessary and sufficient conditions for first order critical points in Appendix A.
We now revisit the central question concerning the choice of functional inequality for dimension reduction. Evidently, the use of the dimensional LSI improves upon the LSI: Theorem 4 shows that sharper error certificates are obtained which are also computationally straightforward to minimize, assuming Conjecture 6 holds. In fact, the following proposition suggests that this is, in some sense, the optimal choice of functional inequality for dimension reduction under the KL divergence. Specifically, minimizing majorant (2.10) is equivalent to minimizing the KL divergence when the target is Gaussian, i.e., the dimensional LSI instrinsically encodes dimension reduction for Gaussian measures. Accordingly, Theorem 4 can be interpreted as the extension of dimension reduction towards non-Gaussian distributions.
Proposition 7.
If $\pi$ is a non-singular Gaussian measure then $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})=\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}))$ for all matrices $U_{r}\in\mathbb{R}^{d\times r}$ with orthonormal columns.
Proof.
Let $\pi=\mathcal{N}(m,C)$ be a Gaussian measure with mean $m$ and covariance $C\succeq 0$. For any orthonormal matrix $U=[U_{r},\,\,U_{\perp}]$ the measure $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ defined as in (2.2) is also Gaussian, but with mean $U_{r}U_{r}^{\top}m$ and covariance $U\mathrm{diag}(U_{r}^{\top}CU_{r},I_{\perp})U^{\top}$. Direct computation shows that
$\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}_{r})=\ln\det(U_{r}^{\top}CU_{r})-\ln\det C-d+r+\operatorname{tr\,}U_{\perp}^{\top}CU_{\perp}+\operatorname{tr\,}U_{\perp}^{\top}mm^{\top}U_{\perp}.$
Combining this with the observation $H(\pi)=\mathbb{E}_{\pi}[C^{-1}(X-m)(X-m)^{\top}C^{-1}]=C^{-1}$ and applying the standard identity $C=M(\pi)-mm^{\top}$ completes the proof. ∎
2.2.1 Numerical Example: Linear Gaussian inverse problems
We consider linear Gaussian inverse problems — applications to a broader class of Bayesian inverse problems are discussed in §2.5. The reference measure $\mu$ is the prior distribution of the model parameters $X$, assumed isotropic Gaussian in $\mathbb{R}^{d}$, and we assume data generated by the model $Y=AX+\epsilon$, where $A\in\mathbb{R}^{m\times d}$ and $\epsilon\sim\mathcal{N}(0,I_{m})$. The posterior measure of $X|Y=y$ is the Gaussian measure $\pi$ with mean $m(\pi)=(I_{d}+A^{\top}A)^{-1}A^{\top}y$ and covariance $C(\pi)=(I_{d}+A^{\top}A)^{-1}$. For any feature $U_{r}\in\mathbb{R}^{d\times r}$, the approximation $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ is also Gaussian with mean $U_{r}U_{r}^{\top}m(\pi)$ and covariance $I_{d}+U_{r}U_{r}^{\top}(C(\pi)-I_{d})U_{r}U_{r}^{\top}$.
Cui et al. [19, Lemma 2.2] considered dimension reduction for Gaussian target measures under the assumption $\mu_{X\mid Y}=\mathbb{E}_{\widetilde{\pi}^{\mathrm{KL}}}[X]$. Provided this holds, they show that the $r$ leading eigenvectors of the relative Fisher information matrix $H$ are globally optimal for (1.2) with the KL divergence. However, since $Y\neq 0$ almost surely, these features are not provably optimal for the typical linear Gaussian inverse problem.
Instead, the globally optimal features can be determined by minimizing $U_{r}\mapsto\mathcal{J}^{\downarrow}(U_{r})$. We evaluate the stability of this objective when the Fisher information matrix and the second moment matrix are empirically estimated from samples. While we use the standard Monte Carlo for $\widehat{H}(\pi||\mu)$ and $\widehat{M}(\pi)$, we use the estimator $\widehat{H}(\pi):=\widehat{H}(\pi||\mu)-\widehat{M}(\pi)+2I_{n_{x}}$ rather than the Monte Carlo estimate of $\mathbb{E}_{\pi}[\nabla\ln\pi^{\otimes 2}]$ as we observed this produced more stable numerical results.
Figure 1 compares the performance of the estimator for a linear Gaussian inverse problem of size $n_{x}=m=50$. The solid black curve depicts the result of minimizing (2.10) with analytical $H(\pi)$ and $M(\pi)$, whereas the solid gray curve depicts the loss corresponding to (2.6). The dashed lines show the result of minimizing (2.10) using empirically estimated matrices of sample size $n_{s}$ — observe that they under predict the true error due to overfitting. Instead, when evaluated on a ‘testing’ set of $\widehat{H}(\pi)$ and $\widehat{M}(\pi)$ computed with $200$ samples, the solid coloured curves show that the estimated $\widehat{U}_{r}^{\downarrow}$ are close to the global minimizer.
Figure 1: Comparison of approximation error certificates for a Gaussian target measure of size $n_{x}=n_{y}=50$. The globally optimally features $U_{r}^{\mathrm{opt}}$ are obtained by minimizing (2.10), and their achieved KL error is shown in black. The certificate obtained from the LSI majorization is shown in gray. The minimizers of the (empirically estimated) dimensional LSI majorant is shown in by the coloured curves, where solid lines correspond to testing error, and dashed lines correspond to training error.
2.2.2 Numerical Example: Rosenbrock
Consider the two-dimensional Rosenbrock distribution $\pi=T_{\#}\mu$ with pushforward $T(z_{1},z_{2})=[z_{1}-0.5,~{}(z_{1}-0.5)^{2}+\sqrt{0.2}z_{2}]$ and standard Gaussian $\mu$. Figure 2 shows the probability density of $\pi$, alongside its approximations $\widetilde{\pi}^{\mathrm{KL}}(U_{1}^{*})$ with feature $U_{1}^{*}$, obtained by minimizing the majorant (2.5), and $\widetilde{\pi}^{\mathrm{KL}}(U_{1}^{\downarrow})$ with feature $U_{1}^{\downarrow}$ obtained by minimizing the majorant (2.10). The computation of each approximation via (2.2) uses Gauss-Hermite quadratures. The relative Fisher information matrix $H(\pi||\mu)$ and moment matrix $M(\pi)$ are similarly computed with quadrature, yielding the Fisher information matrix $H(\pi)$ via (2.14). Note that the tail behaviour of the approximation differs significantly depending on the feature.
Figure 2: (Left) Probability density of target measure. (Centre) Approximate measure using feature $U_{1}^{*}$ determined minimizing (2.5). (Right) Approximate measure using feature $U_{1}^{\downarrow}$ minimizing (2.9).
Each candidate feature $U_{1}\in\mathbb{R}^{2}$ can be parametrized by a unique angle $\theta\in[0,\pi)$ by $U_{1}(\theta)=(\cos(\theta),\sin(\theta))$. For each $\theta$, we compute $\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{1}(\theta)))$, once again using numerical quadratures; this corresponds to the solid blue curve in Figure 3. The shaded gray region corresponds to the majorization and minorization provided in Theorem 2. Minimizing the majorant yields $\theta(U_{1}^{*})$, depicted by the dashed yellow line. In contrast, the shaded blue region correponds to the majorization and minorization from Theorem 4. The minimizer $\theta(U_{1}^{\downarrow})$ of (2.10) is shown by the dashed red line.
We observe that the bounds corresponding to the dimensional LSI are significantly tighter than those corresponding to the (non-dimensional) LSI, consistent with our analysis. Furthermore, we observe the benign non-convex property of $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ as states in Conjecture 6. Surprisingly, we also observe that the dimensional LSI $\theta(U_{1}^{\downarrow})$ minimizer coincides with the global minimizer of the KL divergence, whereas the LSI minimizer $\theta(U_{1}^{*})$ is nearly the worse possible feature.
Figure 3: (Rosenbrock) KL divergence $\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{1}(\theta)))$ with $U_{1}(\theta)=(\cos(\theta),\sin(\theta))$ for $\theta\in[0,\pi]$ (blue line) and its bounds obtain either by the LSI of Theorem 2 (gray shade) or by the dimensional LSI of Theorem 4 (blue shade).
2.3 Choosing the reference measure $\mu$ as the best Gaussian approximation to $\pi$
As a motivation, let us note that for isotropic measure $\pi$ so that $M(\pi)=I_{d}$, the upper bound in (2.9) of Theorem 4 simplifies as
$\mathcal{J}_{\mathrm{KL}}^{\downarrow}(U_{r})$ $=\frac{1}{2}\ln\det(H(\pi))-\frac{1}{2}\ln\det(U_{r}^{\top}H(\pi)U_{r}).$
In addition, if $\pi$ is centered so that $m(\pi)=0$, the lower bound vanishes $\mathcal{J}_{\mathrm{KL}}^{\uparrow}(U_{r})=0$ for any $U_{r}$. In this case, the minimizer of $\mathcal{J}_{\mathrm{KL}}^{\downarrow}(\cdot)$ is known in closed form: it is the matrix $U_{r}^{\downarrow}$ with columns containing the $r$ leading eigenvectors of the Fisher information matrix $H(\pi)$. The bound (2.9) evaluated at this $U_{r}^{\downarrow}$ then simplifies as
$\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}^{\downarrow}))\leq\frac{1}{2}\sum_{k=r+1}^{d}\ln(\lambda_{k}(H(\pi))),$
where $\lambda_{k}(H(\pi))$ is the $k$-th largest eigenvalue of $H(\pi)$. Recalling (2.14), we have $H(\pi)=H(\pi||\gamma)+I_{d}$ so that $U_{r}^{\downarrow}=U_{r}^{*}$, where $U_{r}^{*}$ is the minimizer of the LSI majorant (2.5). This further implies that we have $\lambda_{k}(H(\pi))=1+\lambda_{k}(H(\pi||\mu))$, and using the inequality $\ln(1+t)\leq t$ demonstrates that the above sharpens (2.6).
The same reasoning can be generalized to non-isotropic/non-centered measures $\pi$ by changing the reference measure $\mu$ to be the Gaussian measure $\mu=\mathcal{N}(m(\pi),C(\pi))$. This choice is natural since it is the best Gaussian approximation to $\pi$, in the sense of minimizing $\mu\mapsto\operatorname{D_{\mathrm{KL}}}(\pi||\mu)$ over the set of Gaussian measures. The proof is given in Appendix D.4.
Theorem 8.
Let $\pi$ be a probability measure on $\mathbb{R}^{d}$ with mean $m(\pi)$ and non-singular covariance matrix $C(\pi)$. Consider the Gaussian measure $\mu=\mathcal{N}(m(\pi),C(\pi))$ with same mean and covariance. Then, for any $U_{r}\in\mathbb{R}^{d\times r}$ such that $U_{r}^{\top}C(\pi)U_{r}=I_{r}$, the approximation $\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ as in (2.2) satisfies
$\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(U_{r}))\leq\frac{1}{2}\ln\det(C(\pi)H(\pi))-\frac{1}{2}\ln\det(V_{r}^{\top}H(\pi)V_{r}),$ (2.16)
where $V_{r}=C(\pi)U_{r}$ and where $H(\pi)=\mathbb{E}_{\pi}[\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}x})\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}x})^{\top}]$ is the Fisher information matrix of $\pi$. In particular, the minimizer the above right-hand side is given by $U_{r}^{\downarrow}=C(\pi)^{-1}V_{r}^{\downarrow}$ where $V_{r}^{\downarrow}=[v_{1},\ldots,v_{r}]$ contains the $r$ largest generalized eigenvectors of the matrix pair $(H(\pi),C(\pi)^{-1})$, meaning $H(\pi)v_{k}=\lambda_{k}(H(\pi))C(\pi)^{-1}v_{k},$ with the orthogonality constraint $v_{i}^{\top}C(\pi)^{-1}v_{j}=\delta_{i,j}$. In this case, we obtain
$\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{r}^{\downarrow}))\leq\frac{1}{2}\sum_{k=r+1}^{d}\ln(\lambda_{k}(H(\pi))).$ (2.17)
Theorem 8 suggests that the deviations between $\pi$ and its best Gaussian approximation $\mu=\mathcal{N}(m(\pi),C(\pi))$ are captured by the relative deviations — in the sense of generalized Rayleigh quotients — of the Fisher information matrix $H(\pi)$ and the precision matrix $C(\pi)^{-1}$. Note that for Gaussian target measures $\pi=\mathcal{N}(m,C)$, we have $H(\pi)=C^{-1}$ so that the right-hand side of (2.16) vanishes for any $U_{r}$. This is not surprising because, in this case, $\pi=\mu=\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ for any $U_{r}$.
2.4 Numerical example: Rosenbrock with best Gaussian reference measure $\mu$
We revisit the two-dimensional Rosenbrock distribution from Section 2.2.2 with the reference measure $\mu=\mathcal{N}(m(\pi),C(\pi))$. Figure 4 depicts the density of the optimal Gaussian reference (note it’s anisotropic covariance structure), as well as an approximation $\widetilde{\pi}^{\mathrm{KL}}(U_{1})$ with randomly chosen feature $U_{1}$. Observe that the approximation retains the tail properties from the reference measure.
In Figure 5 we compare the approximation properties when using the standard reference measure $\mu=\mathcal{N}(0,I)$ compared to the optimal Gaussian reference $\mu=\mathcal{N}(m(\pi),C(\pi))$. The red line shows that the optimal reference achieves uniformly lower approximation error than with the standard reference measure shown by the blue line. However, in comparing the performances of the majorants, we note that the majorant in Theorem 8 (shown by the upper red envelope) suggests to select the globally sub-optimal feature direction at $\theta=180$ degrees. In contrast, the minimizer of the majorant from Theorem 4 (upper blue envelope) certifies a lower approximation error, while also coinciding with the globally optimal direction, as previously discussed in Section 2.2.2. It is an interesting research direction to understand when one should or should not construct approximations based on the standard reference measure.
Figure 4: (Left) Probability density of target measure. (Centre) Optimal Gaussian prior with mean $m(\pi)$ and covariance $C(\pi)$. (Right) Approximate measure with feature $U_{r}$ using optimal Gaussian prior.
Figure 5: (Rosenbrock) KL divergence $\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{1}(\theta)))$ with $U_{1}(\theta)=(\cos(\theta),\sin(\theta))$ for $\theta\in[0,\pi]$ (red line) with the optimal Gaussian reference measure $\mu=\mathcal{N}(m(\pi),C(\pi))$ and its bounds obtained by the dimensional LSI of Theorem 8 (red shade). We include for comparison the true approximation error (blue line) and dimensional LSI bounds (blue shade) in Theorem 4 when constructing an approximation using the standard reference measure $\mu=\mathcal{N}(0,1)$.
2.5 Application to Bayesian inverse problems: the data-free setting
For Bayesian inverse problems, the target measure is the posterior $\mathrm{d}\pi^{y}(x)\propto\ell^{y}(x)\mathrm{d}\mu(x)$, where the likelihood function $x\mapsto\ell^{y}(x)$ is proportional to the probability density of observing a given data $y$ knowing $x$. As noticed in [43, 48, 20], in many practical settings the data $y$ only provide information about a few linear features of $x$. Choosing the prior $\mu$ as the reference measure is natural here, and $U_{r}$ can be interpreted as the directions which captures the prior-to-posterior update. The methodology in §2.1 and §2.2 provides an algorithm towards circumventing the impact of high-dimensionality in posterior sampling.
The data-free approach proposed in [21] consists of seeking for low-dimensional features $V_{r}\in\mathbb{R}^{d\times r}$ which are informed by the average realization of data. Formally, these features are characterized by
$\min_{\begin{subarray}{c}V_{r}\in\mathbb{R}^{d\times r}\\ V_{r}^{\mathsf{T}}V_{r}=I_{r}\end{subarray}}\,\mathbb{E}_{Y}\left[\,\min_{\ell_{r}^{Y}:\,\mathbb{R}^{r}\to\mathbb{R}_{+}}\operatorname{D_{\mathrm{KL}}}(\pi^{Y}\,||\,\widetilde{\pi}(V_{r},\ell_{r}^{Y}))\right]\quad\text{for }\mathrm{d}\widetilde{\pi}(x\mid V_{r},\ell_{r}^{y})\propto\ell_{r}^{y}(V_{r}^{\mathsf{T}}x)\mathrm{d}\mu(x).$ (2.18)
In contrast to (1.2), optimality is quantified with respect to the averaged KL divergence with respect to data $Y$, with marginal law given by $\mathrm{d}\rho(y)\propto(\int\ell^{y}(x)\mathrm{d}\mu(x))\mathrm{d}y$. The authors demonstrate an application of the logarithmic Sobolev inequality to derive a majorization to (2.18). We show in Theorem 9 that a tighter majorization can be attained by leveraging the dimensional LSI instead. The proofs of these statements are left to Appendix D.5.
Theorem 9.
Let $\mathrm{d}\pi^{Y}(x)\propto\ell^{Y}(x)\mathrm{d}\mu(x)$ be the posterior distribution of the random vector $X\sim\mu$ conditioned on $Y$, where $\mu=\mathcal{N}(0,I_{d})$ is the standard Gaussian prior and $\ell^{y}(x)$ is the likelihood function of observing $Y=y$ given $X=x$. For any matrix $V_{r}\in\mathbb{R}^{d\times r}$ with $r\leq d$ orthogonal columns, the measure $\mathrm{d}\widetilde{\pi}^{\mathrm{KL},y}(x\mid V_{r})\propto\ell^{\mathrm{KL},y}_{r}(V_{r}^{\top}x)\mathrm{d}\mu(x)$ with $\ell^{\mathrm{KL},y}_{r}(\theta_{r})=\mathbb{E}_{X\sim\mu}[\ell^{y}(X)\mid V_{r}^{\top}X=\theta_{r}]$ satisfies
$\mathbb{E}_{Y}\left[\operatorname{D_{\mathrm{KL}}}(\pi^{Y}||\widetilde{\pi}^{\mathrm{KL},Y}(V_{r}))\right]\leq\frac{1}{2}\ln\det\left(V_{\perp}^{\top}(I_{d}+H_{\mathrm{df}})V_{\perp}\right),$ (2.19)
where $V_{\perp}\in\mathbb{R}^{d\times(d-r)}$ denotes any orthogonal completion to $V_{r}$ and where $H_{\mathrm{df}}\in\mathbb{R}^{d\times d}$ is the data-free diagnostic matrix defined by
$H_{\mathrm{df}}=\mathbb{E}_{X,Y}[\nabla_{x}\ln\ell^{Y}(X)\nabla_{x}\ln\ell^{Y}(X)^{\top}].$ (2.20)
In particular, the minimizer $V_{r}^{\downarrow}$ of the right-hand side of (2.19) is given by $V_{r}^{\downarrow}=[v_{1},\ldots,v_{r}]$, where $v_{k}$ is the $k$-th largest eigenvector of $H_{\mathrm{df}}$. Denoting by $\lambda_{k}(H_{\mathrm{df}})$ the corresponding eigenvalue, (2.19) yields
$\mathbb{E}_{Y}\left[\operatorname{D_{\mathrm{KL}}}(\pi^{Y}\,||\,\widetilde{\pi}^{\mathrm{KL}}\,{}^{Y}(V_{r}))\right]\leq\frac{1}{2}\sum_{k=r+1}^{d}\ln(1+\lambda_{k}(H_{\mathrm{df}})).$ (2.21)
The nomenclature ‘data-free’ refers to the fact that $H_{\mathrm{df}}$ only requires an expectation over the joint law of $(X,Y)$ and does not depend on the observed realization $y$ of the data. Samples from the joint $(X,Y)$ can be easily obtained via Gibb’s algorithm, by first sampling $X\sim\mu$ and then $Y\mid X\sim\pi^{X}$ where $\mathrm{d}\pi^{x}(y)\propto\ell^{y}(x)\mathrm{d}y$.
Compared to the majorant $\mathbb{E}_{Y}[\operatorname{D_{\mathrm{KL}}}(\pi^{Y}||\widetilde{\pi}^{\mathrm{KL},Y}(V_{r}))]\leq\frac{1}{2}\operatorname{tr\,}((I_{d}-V_{r}V_{r}^{\top})H_{\mathrm{df}})$ obtained by [21, §3.3], (2.19) is strictly tighter by $\ln\det(I+A)\leq\operatorname{tr\,}(A)$. While both majorants are minimized by the same feature $V_{r}^{\downarrow}$, the certificate of Cui and Zahm only ensures that
$\mathbb{E}_{Y}\operatorname{D_{\mathrm{KL}}}(\pi\,||\,\widetilde{\pi}^{\mathrm{KL}}(V_{r}^{\downarrow}))\leq\frac{1}{2}\sum_{k=r+1}^{d}\lambda_{k}(H_{\mathrm{df}}),$
whereas Theorem 9 guarantees that the approximation error can be significantly less. In some instances this can be exponentially better, such as for small $r$ when the sum of the trailing eigenvalues is large. Consequently, significantly fewer dimensions $r$ suffice to achieve the same error tolerance than previously considered.
Remark 3.
Proposition 7 extends to the data-free setting, that is, $\mathbb{E}_{Y}\left[\operatorname{D_{\mathrm{KL}}}(\pi^{Y}||\widetilde{\pi}^{\mathrm{KL},Y}(V_{r}))\right]=\frac{1}{2}\ln\det\left(V_{\perp}^{\top}(I_{d}+H_{\mathrm{df}})V_{\perp}\right),$ for any $V_{r}$ whenever $\pi^{Y}$ is Gaussian.
Remark 4.
As noticed in [13], $\mathbb{E}_{Y}\operatorname{D_{\mathrm{KL}}}(\pi^{Y}||\widetilde{\pi}^{\mathrm{KL},Y}(V_{r}))=I(X;Y)-I(V_{r}^{\top}X;Y)$ where $I(X,Y)$ is the mutual information between $X$ and $Y$. Thus, the bound (2.19) yields a lower-bound on the mutual information between $V_{r}^{\top}X$ and $Y$ given by
$I(V_{r}^{\top}X;Y)\geq I(X;Y)-\frac{1}{2}\ln\det\left(V_{\perp}^{\top}(I_{d}+H_{\mathrm{df}})V_{\perp}\right).$
Remark 5.
The reverse dimensional LSI (2.8) results in a lower bound (not shown) involving the data-free lower diagnostic matrix
$G_{\mathrm{df}}=\mathbb{E}_{Y}[\mathbb{E}_{X\mid Y}[\nabla_{x}\ln\ell^{Y}(X)]\mathbb{E}_{X\mid Y}[\nabla_{x}\ln\ell^{Y}(X)]^{\top}].$
This is difficult to compute on account of the nested expectations with respect to the law of $X\mid Y$. A looser bound can be obtained by lower bounding this matrix, in the Loewner sense, by $G^{\prime}_{\mathrm{df}}=\mathbb{E}_{X,Y}[\nabla_{x}\ln\ell^{Y}(X)]\mathbb{E}_{X,Y}[\nabla_{x}\ln\ell^{Y}(X)]^{\top}$. However, $G^{\prime}_{\mathrm{df}}=0$ under the commonly used data likelihood model $Y\mid X\sim\mathcal{N}(g(X),~{}\sigma_{\mathrm{obs}}^{2})$, with parameter-to-observation model $g$, resulting in a vacuous bound.
2.5.1 Applications to Inverse Problems with Generative Modeling Priors
Recent research has focused on Bayesian inverse problems relying on generative modeling priors such as GANs [27]. This adopts an empirical Bayes philosophy wherein the prior is constructed from training samples and is assumed to be supported on a low-dimensional manifold. Mathematically, the prior is codified as the pushforward of a low-dimensional latent variable $Z$ taking values in $\mathbb{R}^{\kappa},~{}\kappa\ll d$, so that $\mu_{X}=\phi_{\#}\mu_{Z}$ for some $\phi:\mathbb{R}^{\kappa}\to\mathbb{R}^{d}$. Without loss of generality, the measure $\mu_{Z}$ can be taken to be standard Gaussian. This procedure significantly reduces the complexity of sampling — it suffices to sample the latent posterior
$\mathrm{d}\pi_{Z}^{y}(z)\propto\ell^{y}(\phi(z))\mathrm{d}\mu_{Z}(z),$ (2.22)
and posterior samples in the model space $X$ can be obtained by the pushforward condition $\pi^{y}_{X}=\phi_{\#}\pi^{y}_{Z}$.
Figure 6: (Left) GAN decoder applied to prior sample $Z\sim\mu_{Z}$. (Right) GAN decoder applied to same prior sample but projected onto features $V_{r}^{\downarrow}$. Increasing $r$ shows which features (in the nominal model space) are informed by data.
We demonstrate that additional dimension reduction can be achieved by accounting for interactions between data, likelihood models, and generative priors. In particular, we consider a geophysical inverse problem in which the nominal parameter space $X$ are discretizations of the subsurface taking values in $3600$ dimensions. Since realizations of the subsurface exhibit regular, repeated layered structures, high-fidelity prior samples can be drawn from a trained GAN prior with only a $300$ dimensional latent space (see Figure 6). Using this GAN prior we compute the error certificate (2.19) and plot it in Figure 7 as a function of rank $r$. This figure shows that data only inform relatively few latent coordinates — e.g., it suffices consider $r=50$ features to obtain an expected KL error of $0.1$ nats. Moreover, domain practitioners can visualize which features are “data-informed” by visualizing the convergence of the GAN output with increasing $r$, as depicted in Figure 6.
Figure 7: The data-free approximation error certificate (2.21) versus rank $r$ for the geophysical inverse problem with generative GAN prior.
3 Dimension reduction with the squared Hellinger distance
It is of interest to consider alternative error metrics in (1.2) to the KL divergence. Notably, [20, 17] considered the squared Hellinger distance $\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\mu)=\frac{1}{2}\int(\sqrt{\mathrm{d}\pi/\mathrm{d}x}-\sqrt{\mathrm{d}\mu/\mathrm{d}x})^{2}\mathrm{d}x$. Straightforward calculation permits us to relate this distance to the variance $\operatorname{Var}_{\mu}(\sqrt{\mathrm{d}\pi/\mathrm{d}\mu})$ following
$\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\mu)=1-\sqrt{1-\operatorname{Var}_{\mu}\left(\sqrt{\frac{\mathrm{d}\pi}{\mathrm{d}\mu}}\right)}.$ (3.1)
As the Poincaré inequality provides gradient-based bounds on variances, it is natural choice to employ it in this context to control the squared Hellinger distance (just as the logarithmic Sobolev inequality naturally relates to the KL divergence). The Poincaré inequality, however, similarly exhibits a “dimension-free” property as the LSI, see e.g. [18, §3], and thus one may wonder whether dimensional refinements can be leveraged to obtain improved majorizations. In this section, we discuss such refinements to the Poincaré inequality for the purpose of dimension reduction in the squared Hellinger distance. Our presentation largely mirrors Section 2.
3.1 Certifiable bounds using the Poincaré Inequality
Let us recall the Poincaré inequality for the standard Gaussian measure $\mu=\mathcal{N}(0,I_{d})$.
Proposition 10 (Theorem 4.7.2. in [8]).
The standard Gaussian measure $\mu=\mathcal{N}(0,I_{d})$ satisfies
$\left\|\int\nabla f(x)\mathrm{d}\mu\right\|_{2}^{2}\leq\operatorname{Var}_{\mu}(f)\leq\int\|\nabla f\|_{2}^{2}\mathrm{d}\mu.$ (3.2)
for any smooth function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We refer to the upper (resp. lower) bound as the (resp. reverse) Poincaré inequality.
By letting $f=\sqrt{\mathrm{d}\pi/\mathrm{d}\mu}$ in (3.2), direct calculation permits to bound the squared Hellinger distance (3.1) from $\pi$ to $\mu=\mathcal{N}(0,I_{d})$ as follow
$1-\sqrt{\left(1-\left\|\mathbb{E}_{\pi}\left[X\sqrt{\frac{\mathrm{d}\mu}{\mathrm{d}\pi}}\right]\right\|^{2}\right)_{+}}\leq\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\mu)\leq 1-\sqrt{\left(1-\frac{\operatorname{tr\,}(H(\pi||\mu)}{4})\right)_{+}},$
where $H(\pi||\mu)=\mathbb{E}_{\pi}[\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})^{\top}]$ is the Fisher information matrix of $\pi$ relative to $\mu$ and $(t)_{+}:=\max(t,0)$. The above left-hand side involves the term $\mathbb{E}_{\pi}[X\sqrt{\frac{\mathrm{d}\mu}{\mathrm{d}\pi}}]$ which, compared to (2.5), is different to the mean of $\pi$. Its computation is, however, difficult when $\pi$ is known up to a normalizing constant which makes it not useable in practice.
With another judicious choice of test function $f$, the same Poincaré inequality yields a tractable upper-bound for the squared Hellinger distance between $\pi$ and a structured approximation $\widetilde{\pi}$ as in (1.1).
Theorem 11 (Adapted from [34]).
Let $\pi$ be a probability measures on $\mathbb{R}^{d}$ and let $\mu=\mathcal{N}(0,I_{d})$ be the standard Gaussian measure. Then, for every matrix $U_{r}\in\mathbb{R}^{d\times r}$ with $r<d$ orthogonal columns, the minimizer of $\ell_{r}\mapsto\operatorname{d_{\mathrm{Hell}}^{2}}(\pi\,||\,\widetilde{\pi}(U_{r},\ell_{r}))$ where $\mathrm{d}\widetilde{\pi}(x\mid U_{r},\ell_{r})\propto\ell_{r}(U_{r}^{\top}x)\mathrm{d}\mu(x)$ is
$\ell^{\mathrm{Hell}}_{r}(\theta_{r}\mid U_{r})=\mathbb{E}_{X\sim\mu}\left[\left.\sqrt{\frac{\mathrm{d}\pi}{\mathrm{d}\mu}}(X)\,\right|\,U_{r}^{\top}X=\theta_{r}\right]^{2}.$ (3.3)
Furthermore, the resulting measure $\widetilde{\pi}^{\mathrm{Hell}}(U_{r})=\widetilde{\pi}(U_{r},\ell^{\mathrm{Hell}}_{r}(U_{r}))$ satisifies
$\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\widetilde{\pi}^{\mathrm{Hell}}(U_{r}))\leq 1-\sqrt{\left(1-\frac{\operatorname{tr\,}(H(\pi||\mu))-\operatorname{tr\,}(U_{r}^{\top}H(\pi||\mu)U_{r})}{4}\right)_{+}},$ (3.4)
where $H(\pi||\mu)=\mathbb{E}_{\pi}[\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})^{\top}]$ is the Fisher information matrix of $\pi$ relative to $\mu$.
The matrix $U_{r}^{*}=[u_{1},\ldots,u_{r}]$ whose columns are the $r$ leading eigenvectors of $H(\pi||\mu)$ is a minimizer of the right-hand side of (3.4), just as with the LSI bound for the KL divergence (c.f. (2.6)). (In fact, [34] show these features certify dimension reduction for all Amari $\alpha$-divergences with $\alpha\in(0,1]$.) This furnishes $\widetilde{\pi}^{\mathrm{Hell}}(U_{r}^{*})$ with the error certificate in the squared Hellinger distance
$\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\widetilde{\pi}^{\mathrm{Hell}}(U_{r}^{*}))\leq 1-\sqrt{\left(1-\frac{1}{4}\sum_{k=r+1}^{d}\lambda_{k}(H(\pi||\mu))\right)_{+}}.$ (3.5)
Whereas the reverse LSI yields the lower bound in Prop. 1, the reverse Poincaré inequality in (3.2) does not lead to a computable minorization for the squared Hellinger distance without explicit reference to $Z_{\pi}$.
3.2 Dimensional Improvements to the Poincaré Inequality
There are several formulations of dimensional Poincaré inequalities for the standard Gaussian $\mu$ in the literature, including:
  • the Bobkov-Ledoux [11] inequality
    $\operatorname{Var}_{\mu}(f)\leq 6\int\|\nabla f\|_{2}^{2}\mathrm{d}\mu-6\int\frac{\|\langle\nabla f,x\rangle\|_{2}^{2}}{d+\|x\|_{2}^{2}}\mathrm{d}\mu,$
  • the Bonnefont-Joulin-Ma inequality [9]
    $\operatorname{Var}_{\mu}(f)\leq\frac{d(d+3)}{d-1}\int\frac{\|\nabla f\|_{2}^{2}}{1+\|x\|_{2}^{2}}\mathrm{d}\mu,$
  • and the first inequality of Bolley-Gentil-Guillin [7], who obtain
    $\operatorname{Var}_{\mu}(f)\leq\int\|\nabla f\|_{2}^{2}\mathrm{d}\mu-\frac{1}{2d}\left(\int(\|x\|_{2}^{2}-d)f\mathrm{d}\mu\right)^{2}$
    by linearizing the dimensional logarithmic Sobolev inequality. [^2] Linearizing refers to the choice of test function $f_{\epsilon}=1+\epsilon g$, where $\int g\mathrm{d}\mu=0$. Substituting into (2.7), scaling $\epsilon\to 0$, and matching terms of order $\epsilon^{2}$ recovers the desired inequality. A similar exercise applied to the LSI (2.4) recovers the Poincaré inequality (3.2).
These inequalities are not mutually comparable as one does not imply the other.
Although we do not explicitly document this, the Bobkov-Ledoux and the Bonnefont-Joulin-Ma inequalities produce worse majorizations than (3.4) obtained from the (non-dimensional) Poincaré inequality. We speculate this is due to their large constants — whereas (3.2) is sharp for $f(x)=x$, it is unclear whether the prefactors of both these inequalities can be further optimized. Furthermore, it can be shown that the Bolley-Gentil-Guillin inequality produces a majorant which explicitly requires the normalizing constant of $\pi$, and is thus of commensurate difficulty to compute when $\pi$ is known up to a constant.
Instead, we rely on another dimensional improvement to the Poincaré inequality proposed by Bolley, Gentil, and Guillin.
Proposition 12 (Theorem 4.3 in [7]).
The variance of a smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ with respect to the standard Gaussian $\mu=\mathcal{N}(0,I_{d})$ satisfies the dimensional Poincaré inequality
$\operatorname{Var}_{\mu}(f)\leq\int\|\nabla f\|_{2}^{2}\mathrm{d}\mu-\int\frac{\|f-\int f\mathrm{d}\mu-\langle\nabla f,x\rangle\|_{2}^{2}}{d+\|x\|_{2}^{2}}\mathrm{d}\mu.$ (3.6)
Remark 6.
We are not aware of a reverse dimensional Poincaré inequality. The proof technique used in [8, Thm 6.7.3] to derive the reverse dimensional LSI relies on an algebraic identity specific to the LSI entropy function $x\mapsto x\ln x$, which cannot be adapted here. Alternatively, linearizing the reverse dimensional LSI merely recovers the (non-dimensional) reverse Poincaré inequality in Proposition 10.
Evidently (3.6) improves on the Poincaré inequality, although the difference vanishes in the limit as $d\to\infty$. Note that the dimensional Poincaré is also tight for the test function $f(x)=x$.
We apply the dimensional Poincaré inequality to obtain a majorization for the squared Hellinger approximation error. The proof can be found in Appendix D.6.
Theorem 13.
Under the same setting as Theorem 11, we have
$\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\,\widetilde{\pi}^{\mathrm{Hell}}(U_{r}))\leq 1-\sqrt{\left(1-\frac{\operatorname{tr\,}(H(\pi||\mu))-\operatorname{tr\,}(U_{r}^{\top}H(\pi||\mu)U_{r})}{4}+\delta_{r}(U_{r},\,\operatorname{d_{\mathrm{Hell}}^{2}}(\pi,\,\widetilde{\pi}^{\mathrm{Hell}}(U_{r})))\right)_{+}}$ (3.7)
with
$\delta_{r}(U_{r},y)=\frac{(1-(1-y)^{2}-\frac{1}{2}\operatorname{tr\,}(M(\pi))+\frac{1}{2}\operatorname{tr\,}(U_{r}^{\top}M(\pi)U_{r})+\frac{1}{2}(d-r))^{2}}{\operatorname{tr\,}(M(\pi))-\operatorname{tr\,}(U_{r}^{\top}M(\pi)U_{r})+d-r},$ (3.8)
where $H(\pi||\mu)=\mathbb{E}_{\pi}[\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})\nabla\ln(\frac{\mathrm{d}\pi}{\mathrm{d}\mu})^{\top}]$ is the Fisher information matrix of $\pi$ relative to $\mu$ and $M(\pi)=\mathbb{E}_{\pi}[XX^{\top}]$ the second moment matrix.
Remark 7.
For isotropic measure $\pi$ such that $M(\pi)=I_{d}$ we obtain
$\epsilon_{r}(U_{r},y)=\frac{(1-(1-y)^{2}))^{2}}{2(d-r)}.$
Observe that the function $\mathbb{R}^{d\times r}\times\mathbb{R}\ni(U_{r},y)\mapsto\delta_{r}(U_{r},y)\in\mathbb{R}$ defined in (3.8) is strictly positive — the numerator is evidently positive, while the denominator is positive since $M(\pi)\succ 0$. By inspection, this implies that the majorant (3.7) improves upon the majorant (3.4). Unfortunately, the squared Hellinger loss appears in both the left and right hand side of (3.7), and is therefore difficult to optimize to determine linear features $U_{r}$. Instead, one might hope to evaluate (3.7) with the features $U_{r}^{*}$ minimizing (3.4) to obtain a tighter certificate for the approximation error. Surprisingly, to do so one cannot simply bootstrap the upper bound from (3.4) to further bound the right hand side of (3.7)— instead, one requires a lower bound on the approximation error in order to leverage the improvements from (3.7). Although non-vacuous lower bounds can be obtained from applying the reverse Poincaré inequality (3.2), we do not expect such bounds to be easily computable. It therefore remains an open question as to how we can make use of the improvements brought on by the dimensional Poincaré inequality, and we leave exploring this direction to future work.
Appendix A Necessary and sufficient conditions for first order optimality of (2.10)
Encoding the orthogonality constraint into (2.10) with Lagrange multiplier $S=S^{\top}\in\mathbb{R}^{r\times r}$ and ignoring constant terms yields the Lagrangian
$L(U_{r},S)=\frac{\operatorname{tr\,}\left(U_{r}^{\top}M(\pi)U_{r}\right)+\ln\det\left(U_{r}^{\top}H(\pi)^{-1}U_{r}\right)+\operatorname{tr\,}\left(S(I_{r}-U_{r}^{\top}U_{r})\right)}{2}.$
First-order optimality asserts that
$\nabla_{U_{r}}L(U_{r},S)=M(\pi)U_{r}+H(\pi)^{-1}U_{r}\left(U_{r}^{\top}H(\pi)^{-1}U_{r}\right)^{-1}-U_{r}S=0$ (A.1)
for all orthonormal $U_{r}$. Left multiplying (A.1) by $U_{r}^{\top}$ thus recovers $S=U_{r}^{\top}M(\pi)U_{r}+I_{d},$ and (A.1) is equivalently
$U_{r}=(I_{d}-U_{r}U_{r}^{\top})M(\pi)U_{r}+H(\pi)^{-1}U_{r}\left(U_{r}^{\top}H(\pi)^{-1}U_{r}\right)^{-1}.$ (A.2)
Re-arranging (A.2) also gives the alternative fixed point condition
$U_{r}=H(\pi)\left(I_{d}-(I_{d}-U_{r}U_{r}^{\top})M(\pi)\right)U_{r}(U_{r}^{\top}H(\pi)^{-1}U_{r}).$ (A.3)
We are not aware of any closed form solutions to (A.2) or (A.3).
Observe that left multiplying (A.3) by primal feasible $U_{r}^{\top}$ and $U_{\perp}^{\top}$ recovers
$U_{r}^{\top}\left(H(\pi)-H(\pi)M(\pi)+H(\pi)U_{r}U_{r}^{\top}M(\pi)\right)U_{r}(U_{r}^{\top}H(\pi)^{-1}U_{r})$ $=I_{r}$ (A.4)
$U_{\perp}^{\top}\left(H(\pi)-H(\pi)M(\pi)+H(\pi)U_{r}U_{r}^{\top}M(\pi)\right)U_{r}$ $=0_{d-r,r}$ (A.5)
However, when $r>1$, (A.4) implies that $U_{r}^{\top}\left(H(\pi)-H(\pi)M(\pi)+H(\pi)U_{r}U_{r}^{\top}M(\pi)\right)U_{r}$ is symmetric, as it is the inverse of a symmetric matrix. This means
$U_{r}^{\top}(H(\pi)M(\pi)-M(\pi)H(\pi))U_{r}=(U_{r}^{\top}H(\pi)U_{r})(U_{r}^{\top}M(\pi)U_{r})-(U_{r}^{\top}M(\pi)U_{r})(U_{r}^{\top}H(\pi)U_{r}),$
or, expanding $H(\pi)$ as (2.14) and condensing notation via the matrix commutator $[A,B]=AB-BA$, we have that
$U_{r}^{\top}[H(\pi||\mu),M(\pi)]\,U_{r}$ $=[U_{r}^{\top}H(\pi||\mu)U_{r},\,U_{r}^{\top}M(\pi)U_{r}]$ (A.6)
$U_{\perp}^{\top}(H(\pi||\mu)+M(\pi)+H(\pi||\mu)M(\pi)-M(\pi)^{2})U_{r}$ $=U_{\perp}^{\top}(H-M(\pi))U_{r}U_{r}^{\top}M(\pi)U_{r}.$ (A.7)
To reiterate: $U_{r}$ is a critical point of (2.10) if and only if it satisfies (A.6) and (A.7). While (A.6) is satisfied if $U_{r}$ contains any $r$ eigenvectors of $H(\pi||\mu)$ or $M(\pi)$, these are in general these are not the only solutions. We examine both of these cases individually:
  1. 1.
    suppose $U_{r}$ contains any $r$ eigenvectors of the second moment matrix $M(\pi)$, so that $M(\pi)U_{r}=U_{r}S$ where $S=\mathrm{diag}(s_{1},\ldots,s_{r})$. Then (A.7) becomes
    $U_{\perp}^{\top}H(\pi||\mu)U_{r}=0_{d-r,r}\,.$
    This condition cannot be satisfied as vectors cannot be simultaneously orthogonal, $M(\pi)$-orthogonal, and $H(\pi||\mu)$-orthogonal, unless any pairs of these matrices commute.
  2. 2.
    suppose $U_{r}$ contains any $r$ eigenvectors of $H(\pi||\mu)$ with $H(\pi||\mu)U_{\perp}=U_{\perp}T_{\perp}$. Then (A.7) becomes
    $(I_{\perp}+T_{\perp})U_{\perp}^{\top}M(\pi)U_{r}-U_{\perp}^{\top}M(\pi)^{2}U_{r}=-(U_{\perp}^{\top}M(\pi)U_{r})(U_{r}^{\top}M(\pi)U_{r}).$
    Numerically, we observe for Gaussian target distributions, that this condition is not met. In other words, the leading eigenvectors of the relative Fisher information matrix, while globally optimal for (2.5), are not even local minima of (2.10) in general.
    Note that this does not contradict Lemma 2.2 in [43]. In their setting, they assume the mean of $\pi^{\mathrm{KL}}_{r}$ matches the mean of $\pi$ — in general, this condition is not true. However, when this criteria is met, then it can be shown that (A.7) is also satisfied.
These results suggest that numerical optimization is required in general to obtain critical points of (2.10).
Remark 8.
Observe that (A.4) allows the component $(U_{r}^{\top}H(\pi)^{-1}U_{r})$ appearing in (2.10) to be replaced by a function involving only $H(\pi)$ and $M$ at optimality, which does not require computing matrix inverses. This may offer computational speedups, but we do not investigate this further in the current work.
Appendix B Auxiliary Matrix Determinant Identity
Sylvester’s determinant theorem allows us to write
$\det\left(U_{\perp}^{\top}H(\pi)U_{\perp}\right)$ $=\det\left(U_{\perp}^{\top}(H(\pi)-I_{d})U_{\perp}+I_{\perp}\right)=\det\left((H(\pi)-I_{d})(U_{\perp}U_{\perp}^{\top})+I_{d}\right)$
$=\det\left((H(\pi)-I_{d})(I_{d}-U_{r}U_{r}^{\top})+I_{d}\right)=\det\left(H(\pi)-H(\pi)\,U_{r}U_{r}^{\top}-(I_{d}-U_{r}U_{r}^{\top})+I_{d}\right)$
$=\det\left(H(\pi)-H(\pi)\,U_{r}U_{r}^{\top}+U_{r}U_{r}^{\top}\right)=\det\left(H(\pi)\right)\det\left(I_{d}-U_{r}U_{r}^{\top}+H(\pi)^{-1}U_{r}U_{r}^{\top}\right)$
$=\det\left(H(\pi)\right)\det\left(I_{d}+(H(\pi)^{-1}U_{r}-U_{r})U_{r}^{\top}\right)=\det\left(H(\pi)\right)\det\left(I_{r}+U_{r}^{\top}(H(\pi)^{-1}U_{r}-U_{r})\right)$
$=\det\left(H(\pi)\right)\det\left(U_{r}^{\top}H(\pi)^{-1}U_{r}\right),$
assuming $H(\pi)$ is invertible.
Appendix C Derivation of Functional Inequalities
C.1 Gaussian Logarithmic Sobolev Inequality
Numerous derivations of the Gaussian logarithmic Sobolev inequality are known, but we prefer the presentation of [8] as this also conveniently encapsulates the reverse LSI. The following is a condensed version of [8, Thm 5.5.2], which is sufficient to obtain Theorem 1.
Theorem 14 (Local Logarithmic Sobolev Inequality).
A Markov diffusion process $\{X_{t},~{}t\geq 0\}$ satisfies a curvature dimension inequality $\textsc{CD}(0,\infty)$ if and only if it satisfies the local logarithmic Sobolev inequality
$P_{t}(f\ln f)-P_{t}f\ln P_{t}f\leq tP_{t}\left(\frac{\Gamma f}{f}\right)$ (C.1)
and the reverse local logarithmic Sobolev inequality
$P_{t}(f\ln f)-P_{t}f\ln P_{t}f\geq t\,\frac{\Gamma(P_{t}f)}{P_{t}f}$ (C.2)
for smooth positive functions $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$.
We emphasize that the presentation as local logarithmic Sobolev inequalities, where local refers to the law of the diffusion processes at time $t<\infty$, is key to obtaining the reverse inequality. On the contrary, it is customary to present the ergodic ($t\to\infty$) formulation of the logarithmic Sobolev inequality, but at the expense of obtaining a vacuous reverse inequality.
The central idea towards converting Eqs (C.1) and (C.2) into Theorem 1 is to specialize to the $d$-dimensional heat process
$\mathrm{d}X_{t}=\mathrm{d}W_{t},\quad X_{0}=x_{0},$ (C.3)
where $W_{t}$ denotes $d$-dimensional Brownian motion. While the heat process does not have an invariant measure, it suffices to observe that the law of $X_{t}$ at any time $t>0$ is isotropic Gaussian with mean $x_{0}$ and marginal variance $t$. It is straightforward to verify that this Markov diffusion process satisfies the $\mathrm{CD}(0,d)$ curvature dimension inequality333also synonymously referred to as the Bakry-Émery criterion.; see [8, Definition 1.16.1] for precise details. In addition, note that if a diffusion process satisfies $\textsc{CD}(\rho,n)$, then it also satisfies $\textsc{CD}(\rho^{\prime},n^{\prime})$ for any $n^{\prime}\geq n$ and $\rho^{\prime}\leq\rho$, hence the heat process is a valid choice for the local logarithmic Sobolev inequality.
The Markov semigroup operator of the heat process is given by $P_{t}f(x):=\mathbb{E}[f(X_{t})\mid X_{0}=x]=\mathbb{E}_{Z\sim\mathcal{N}(0,I_{d})}[f(x+tZ)],$ while its carré du champ operator is
$\Gamma f:=\Gamma(f,f)=\frac{1}{2}\|\nabla f\|_{2}^{2}.$ (C.4)
Accordingly, fix $t=1$ and, without loss of generality, consider the initial condition $x_{0}=0$ (alternatively, we can choose the test functions $g(\cdot)=f(\cdot-x_{0})$). Then, the law of $X_{1}|X_{0}=x_{0}$ is the $d$ dimensional isotropic Gaussian $\mu$, and we have $\textsc{Ent}_{\mu}(f)=\textsc{Ent}_{P_{1}}(f)$, which matches the left hand side of both (C.1) and (C.2). Substituting the carré du champ (C.4) into (C.1) and simplifying via the logarithmic derivative obtains the desired upper bound (2.4). To obtain the reverse inequality, note $P_{1}f(x_{0})=\mathbb{E}_{Z\sim\mu}[f(x_{0}+Z)]$ so that $\nabla_{x}P_{1}f(x_{0})=\mathbb{E}_{Z\sim\mu}[\nabla_{z}f(x_{0}+Z)]$. Evaluating the carré du champ $\Gamma P_{1}f(0)=\frac{1}{2}\|\mu(\nabla f)\|_{2}^{2}$ obtains (C.2) and concludes the proof.
C.2 Dimensional Gaussian Logarithmic Sobolev Inequality
A proof of the dimensional logarithmic Sobolev inequality (and its reverse) for general Markov diffusion processes satisfying specific curvature-dimension inequalities is provided in [8, Thm 6.7.4]. Specifying to the heat process, as above, then recovers the (weaker) dimensional Gaussian LSI of [10].
However, instead we choose to follow the original presentation of [10], later improved by Dembo [22], as this approach clarifies the dependence on the Gaussianity of the reference measure. Neither reference documents the reverse inequality, but these authors were undoubtedly aware of its existence. Since we were unable to find this result explicitly in the literature, we provide its statement and proof below.
The proof proceeds by transforming the Gaussian LSI into Euclidean form by choosing the test function $f(x)=((2\pi)^{-d/2}\exp(-\frac{1}{2}\|x\|_{2}^{2}))^{-1}g(x)$, with $g:\mathbb{R}^{d}\to\mathbb{R}_{+}$ normalized as $\int g(x)\mathrm{d}x=1$ (so that $\int f\mathrm{d}\mu=1$). Applied to the inequalities in (2.4), this yields the Euclidean logarithmic Sobolev inequality
$\int g(x)\ln g(x)\mathrm{d}x\leq\frac{1}{2}\int\frac{\|\nabla g(x)\|_{2}^{2}}{g(x)}\mathrm{d}x-\frac{d}{2}\ln(2\pi)-d,$ (C.5)
as well as the reverse Euclidean logarithmic Sobolev inequality
$\int g(x)\ln g(x)\mathrm{d}x\geq\frac{1}{2}\left\|\int xg(x)\mathrm{d}x\right\|_{2}^{2}-\frac{d}{2}\ln(2\pi)-\frac{1}{2}\int g\|x\|_{2}^{2}\mathrm{d}x,$ (C.6)
after noting that $\int\partial_{k}g(x)\mathrm{d}x=0$ for all $1\leq k\leq d$.
The goal is to bootstrap this to obtain dimensional improvements to the Euclidean LSI. The original insight of [10] is to recognize that the test functions $g$ are probability density functions. This suggests considering $g(x)=(T^{-1})_{\#}f(x)=|\det\nabla T(x)|f(T(x))$ for $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$ and $\int f(x)\mathrm{d}x=1$, i.e., the pushforward of a probability density $f$ under diffeomorphisms $T^{-1}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and optimizing the bounds with respect to $T$ (or $T^{-1}$). Specifically, Bakry and Ledoux considered the map $T(x)=\sigma x$, optimizing over the scalar parameter $\sigma>0$. Dembo [22] generalized this construction to maps $T(x)=Ax$ with symmetric positive definite matrices $A=A^{\top}\succ 0$.
Applying this choice of test function into (C.5), we obtain
$\int f(x)\ln f(x)\mathrm{d}x\leq\frac{1}{2}\operatorname{tr\,}\left(A^{2}\left(\int\frac{\nabla f(x)^{\otimes 2}}{f(x)}\mathrm{d}x\right)\right)-\frac{d}{2}\ln(2\pi)-d-\frac{1}{2}\ln\det(A^{2}),$
and straightforward computation shows that $A^{-2}=\int(\nabla\ln f(x))^{\otimes 2}f(x)\mathrm{d}x$ optimizes the bound. This yields the dimensional Euclidean LSI
$\int f(x)\ln f(x)\mathrm{d}x\leq\frac{1}{2}\ln\det\left(\int\frac{\nabla f(x)^{\otimes 2}}{f(x)}\mathrm{d}x\right)-\frac{d}{2}\ln(2\pi e).$ (C.7)
The same exercise applied to (C.6) yields
$\int f(x)\ln f(x)\mathrm{d}x\geq\frac{1}{2}\left\|A^{-1}\int xf(x)\mathrm{d}x\right\|_{2}^{2}-\frac{d}{2}\ln(2\pi)-\frac{1}{2}\int\|A^{-1}x\|_{2}^{2}f(x)\mathrm{d}x-\ln\det A.$
For this lower bound, the optimal matrix is $A^{2}=\int x^{\otimes 2}f(x)\mathrm{d}x-\left(\int xf(x)\mathrm{d}x\right)^{\otimes 2}$, from which we obtain the dimensional reverse Euclidean LSI
$\int f(x)\ln f(x)\mathrm{d}x\geq-\frac{1}{2}\ln\det\left(\int x^{\otimes 2}f(x)\mathrm{d}x-\left(\int xf(x)\mathrm{d}x\right)^{\otimes 2}\right)-\frac{d}{2}\ln(2\pi e).$ (C.8)
It remains to convert these inequalities for the Gaussian measure by considering test functions $f(x)=g(x)(2\pi)^{d/2}\exp(-\frac{1}{2}\|x\|_{2}^{2})$, normalized $\int f(x)\mathrm{d}x=1$ so that $\int g\mathrm{d}\mu=1$. Substituting this into (C.7) obtains (2.7), and (C.8) obtains (2.8), by applying the Gaussian integration by parts identity $\int x_{k}g(x)\mathrm{d}\mu(x)=\int\partial_{k}g(x)\mathrm{d}\mu(x)$.
Appendix D Miscellaneous Proofs
D.1 Proof of Lower Bound in Theorem 2
We only detail the proof of the lower bound in Theorem 2 below since the upper bound is extensively documented in [48].
Proof.
Without loss of generality, we assume that $\ell:\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ is normalized such that $\int\ell\mathrm{d}\mu=1$. Given a matrix $U_{r}\in\mathbb{R}^{d\times r}$ with orthogonal columns and $x_{r}\in\mathbb{R}^{r}$, we consider the function $f:\mathbb{R}^{d-r}\rightarrow\mathbb{R}$ defined by
$f(x_{\perp})=\frac{\ell(U_{r}x_{r}+U_{\perp}x_{\perp})}{\widetilde{\ell}^{\mathrm{KL}}_{r}(x_{r})},$ (D.1)
where ~$\widetilde{\ell}^{\mathrm{KL}}_{r}$ as in (2.1) and $U_{\perp}$ is any orthogonal complement of $U_{r}$. By definition of ~$\widetilde{\ell}^{\mathrm{KL}}_{r}$, we have $\int f\mathrm{d}\mu_{\perp}=1$. Thus, the LSI (2.4) for the $(d-r)$ dimensional Gaussian measure $\mu_{\perp}$ yields
$\frac{1}{2}\left\|\int\nabla f\mathrm{d}\mu_{\perp}\right\|^{2}\leq\int f\ln f\mathrm{d}\mu_{\perp}.$
Next, we integrate of the above inequality with respect to $x_{r}$ against the probability measure $\mathrm{d}\pi_{r}(x_{r})=\widetilde{\ell}^{\mathrm{KL}}_{r}(x_{r})\mathrm{d}\mu_{r}(x_{r})$, where $\mu_{r}$ is the $r$-dimensional isotropic Gaussian measure (recall that $\int\ell^{\mathrm{KL}}_{r}\mathrm{d}\mu_{r}=\int\ell\mathrm{d}\mu=1$). The right-hand side becomes
$\int\left(\int f\ln f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}$ $=\int\ell(x)\ln\frac{\ell(x)}{\widetilde{\ell}^{\mathrm{KL}}_{r}(U_{r}^{\top}x)}\mathrm{d}\mu(x)=\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{r})).$ (D.2)
Using Jensen’s inequality, the left-hand side satisfies
$\int\left(\frac{1}{2}\left\|\int\nabla f\mathrm{d}\mu_{\perp}\right\|^{2}\right)\mathrm{d}\pi_{r}$ $\geq\frac{1}{2}\left\|\int\left(\int\nabla f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}\right\|^{2}=\frac{1}{2}\left\|U_{\perp}^{\top}\int\nabla\ln\ell\mathrm{d}\pi\right\|^{2}.$
It remains to show that
$\int\nabla\ln\ell\mathrm{d}\pi=\int\nabla\ell\mathrm{d}\mu=-\int\ell\nabla\ln\left(\frac{\mathrm{d}\mu}{\mathrm{d}x}\right)\mathrm{d}\mu=\int x\ell\mathrm{d}\mu=\mathbb{E}_{\pi}[X],$
where we have used integration by parts. Combining the above relations yields (2.5). ∎
D.2 Proof of Theorem 4
Proof.
As in the proof of Theorem 1, we consider the function $f:\mathbb{R}^{d-r}\rightarrow\mathbb{R}$ defined as in (D.1) with $\ell$ assumed to be normalized such that $\int\ell\mathrm{d}\mu=1$. Applying the dimensional LSI (2.7) with $\mu_{\perp}$ yields
$\int f\ln f\mathrm{d}\mu_{\perp}$ $\leq\frac{1}{2}\int|x_{\perp}|^{2}f\mathrm{d}\mu_{\perp}-\frac{d-r}{2}+\frac{1}{2}\ln\det\left(\int\left(\nabla\ln f-x_{\perp}\right)^{\otimes 2}f\mathrm{d}\mu_{\perp}\right)$ (D.3)
Next, we integrate the inequality above with respect to $x_{r}$ against the probability measure $\mathrm{d}\pi_{r}=\widetilde{\ell}^{\mathrm{KL}}_{r}\mathrm{d}\mu_{r}$, where $\mu_{r}$ is the $r$-dimensional isotropic Gaussian measure. As before, following (D.2), the left-hand side of (D.3) becomes $\int\left(\int f\ln f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}=\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{r}))$. The first term in the right-hand of (D.3) side becomes
$\int\left(\frac{1}{2}\int|x_{\perp}|^{2}f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}$ $=\frac{1}{2}\int|U_{\perp}^{\top}x|^{2}\ell\mathrm{d}\mu=\frac{1}{2}\int|U_{\perp}^{\top}x|^{2}\mathrm{d}\pi.$
Finally, using Jensen’s inequality (noting that $S\mapsto\ln\det(S)$ is concave), the last term in the right-hand of (D.3) side satisfies
$\int\frac{1}{2}\ln\det\left(\int\left(\nabla\ln f-x_{\perp}\right)^{\otimes 2}f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}$ $\leq\frac{1}{2}\ln\det\left(\int\left(\int\left(\nabla\ln f-x_{\perp}\right)^{\otimes 2}f\mathrm{d}\mu_{\perp}\right)\mathrm{d}\pi_{r}\right)$
$=\frac{1}{2}\ln\det\left(\int\left(U_{\perp}^{\top}(\nabla\ln\ell-x)\right)^{\otimes 2}\mathrm{d}\pi\right).$
Injecting the three above relations in (D.3) yields the right-bound of (2.9). The lower bound is obtained in a similar way. This concludes the proof. ∎
D.3 Proof of Proposition 5
We require the following lemma.
Lemma 15.
Let $v\in\mathbb{R}^{d}$ such that $v^{\top}v>0$ and $W\in\mathbb{R}^{d\times k}$, $1\leq k<d$, such that $W^{\top}W\succeq 0$. Then we have the inequality
$\ln\det v^{\top}v+\ln\det W^{\top}W\leq\ln\det\begin{pmatrix}v^{\top}v&v^{\top}W\\ W^{\top}v&W^{\top}v\end{pmatrix}.$ (D.4)
Proof.
Re-write (D.4) as
$\ln\det\begin{pmatrix}v^{\top}v&\\ &I_{k}\end{pmatrix}+\ln\det\begin{pmatrix}1&\\ &W^{\top}W\end{pmatrix}\leq\ln\det\begin{pmatrix}v^{\top}v&v^{\top}W\\ W^{\top}v&W^{\top}v\end{pmatrix},$
where $I_{k}$ denotes the $k\times k$ identity matrix. Re-arranging then gives
$\ln\det\begin{pmatrix}1&\theta^{\top}\\ \theta&I_{k}\end{pmatrix}=\ln\det\left(I_{k+1}+\begin{pmatrix}0\\ \theta\end{pmatrix}\begin{pmatrix}0&\theta^{\top}\end{pmatrix}\right)\geq 0$
with $\theta=\frac{1}{\sqrt{v^{\top}v}}(W^{\top}W)^{-\frac{1}{2}}W^{\top}v$. But since $\det\left(I_{k+1}+\begin{pmatrix}0\\ \theta\end{pmatrix}\begin{pmatrix}0&\theta^{\top}\end{pmatrix}\right)=1+\|\theta\|_{2}^{2}$, this concludes the proof. ∎
We now demonstrate that $U_{r}\mapsto\mathcal{J}^{\downarrow}_{d,r}(U_{r})$ is monotonically decreasing for nested sequences.
Proof.
Denoting $U_{r}=\begin{pmatrix}u_{1}&\ldots u_{r}\end{pmatrix}$ and $U_{\perp,r}=\begin{pmatrix}u_{r+1}&\ldots&u_{d}\end{pmatrix}\in\mathbb{R}^{d\times(d-r)}$ for $1\leq r<d$, then
$\mathcal{J}^{\downarrow}(U_{r})=\frac{1}{2}\operatorname{tr\,}U_{\perp,r}^{\top}M(\pi)U_{\perp,r}-\frac{d-r}{2}+\frac{1}{2}\ln\det\left(U_{\perp,r}^{\top}H(\pi)U_{\perp,r}\right)$ (D.5)
and
$\mathcal{J}^{\downarrow}(U_{r+1})=\frac{1}{2}\operatorname{tr\,}U_{\perp,r+1}^{\top}M(\pi)U_{\perp,r+1}-\frac{d-r-1}{2}+\frac{1}{2}\ln\det\left(U_{\perp,r+1}^{\top}H(\pi)U_{\perp,r+1}\right).$
Demonstrating $\mathcal{J}^{\downarrow}(U_{r+1})\leq\mathcal{J}^{\downarrow}(U_{r})$ amounts to the inequality
$1-u_{r+1}^{\top}M(\pi)u_{r+1}+\ln\det\left(U_{\perp,r+1}^{\top}H(\pi)U_{\perp,r+1}\right)\leq\ln\det\left(U_{\perp,r}^{\top}H(\pi)U_{\perp,r}\right).$
The key idea is to recognize that by defining the orthonormal sub-matrix
$V=\begin{pmatrix}u_{1}&\ldots&u_{r}&u_{r+2}&\ldots&u_{d}\end{pmatrix}\in\mathbb{R}^{d\times(d-1)},$
we have
$0\leq\mathcal{J}^{\downarrow}(V)=\frac{u_{r+1}^{\top}M(\pi)u_{r+1}-1}{2}+\frac{1}{2}\ln\det\left(u_{r+1}^{\top}H(\pi)u_{r+1}\right),$
where non-negativity is ensured since
$\operatorname{D_{\mathrm{KL}}}(\pi\,\|\,\widetilde{\pi}^{\mathrm{KL}}(U^{\downarrow}_{d-1}))\leq\min_{U_{d-1}}\mathcal{J}^{\downarrow}(U_{d-1})\leq\mathcal{J}^{\downarrow}(V).$
It therefore suffices to show that
$\ln\det\left(u_{r+1}^{\top}H(\pi)u_{r+1}\right)+\ln\det\left(U_{\perp,r+1}^{\top}H(\pi)U_{\perp,r+1}\right)\leq\ln\det\left(U_{\perp,r}^{\top}H(\pi)U_{\perp,r}\right).$
But this follows from applying Lemma 15 with $v=H(\pi)^{-\frac{1}{2}}u_{r+1}$ and $W=H(\pi)^{-\frac{1}{2}}U_{\perp,r+1}$. ∎
D.4 Proof of Theorem 8
Proof.
Let $T(x)=C(\pi)^{-1/2}(x-m(\pi))$ be the affine map such that $\overline{\pi}=T_{\sharp}\pi$ is centered and isotropic, that is, $m(\overline{\pi})=0$ and $M(\overline{\pi})=I_{d}$. Applying Theorem 4 to $\overline{\pi}$ yields
$\operatorname{D_{\mathrm{KL}}}(\overline{\pi}||\overline{\widetilde{\pi}^{\mathrm{KL}}}(\overline{U_{r}}))\leq\frac{1}{2}\ln\det(H(\overline{\pi}))-\frac{1}{2}\ln\det(\overline{U_{r}}^{\top}H(\overline{\pi})\overline{U_{r}}),$ (D.6)
for any matrix $\overline{U_{r}}\in\mathbb{R}^{d\times r}$ with orthogonal columns. Here $\overline{\widetilde{\pi}^{\mathrm{KL}}}(\overline{U_{r}})=\overline{\widetilde{\pi}^{\mathrm{KL}}}(\cdot\mid\overline{U_{r}})$ is defined by $\mathrm{d}\overline{\widetilde{\pi}^{\mathrm{KL}}}(\overline{x}\mid\overline{U_{r}})=\overline{\ell^{\mathrm{KL}}_{r}}(\overline{U_{r}}^{\top}\overline{x})\mathrm{d}\overline{\mu}(\overline{x}),$ where $\overline{\mu}=\mathcal{N}(0,I_{d})$ and $\overline{\ell^{\mathrm{KL}}_{r}}(\theta_{r})=\mathbb{E}_{\overline{X}\sim\overline{\mu}}[\frac{\mathrm{d}\overline{\pi}}{\mathrm{d}\overline{\mu}}(\overline{X})\mid\overline{U_{r}}^{\top}\overline{X}=\theta_{r}]$. Next, we use a change of variable $x=T^{-1}(\overline{x})$ to show that (D.6) yields (2.16). We can write
$\overline{\ell^{\mathrm{KL}}_{r}}(\overline{U_{r}}^{\top}\overline{x})$ $=\mathbb{E}_{\overline{X}\sim\overline{\mu}}\left[\left.\frac{\mathrm{d}\overline{\pi}}{\mathrm{d}\overline{\mu}}(\overline{X})\right|\overline{U_{r}}^{\top}\overline{X}=\overline{U_{r}}^{\top}\overline{x}\right]$
$=\mathbb{E}_{X\sim\mu}\left[\left.\frac{\mathrm{d}\pi}{\mathrm{d}\mu}(X)\right|\overline{U_{r}}^{\top}T(X)=\overline{U_{r}}^{\top}T(x)\right]$
$=\mathbb{E}_{X\sim\mu}\left[\left.\frac{\mathrm{d}\pi}{\mathrm{d}\mu}(X)\right|\overline{U_{r}}^{\top}C(\pi)^{-1/2}X=\overline{U_{r}}^{\top}C(\pi)^{-1/2}x\right]=\ell^{\mathrm{KL}}_{r}(U_{r}^{\top}x),$
where
$U_{r}=C(\pi)^{-1/2}\overline{U_{r}}.$
We deduce that $\mathrm{d}\overline{\widetilde{\pi}^{\mathrm{KL}}}(\overline{U_{r}})=\ell^{\mathrm{KL}}_{r}(U_{r}^{\top}T^{-1}(\cdot))\mathrm{d}(T_{\sharp}\mu)=T_{\sharp}\widetilde{\pi}^{\mathrm{KL}}(U_{r})$ so that
$\operatorname{D_{\mathrm{KL}}}(\overline{\pi}||\overline{\widetilde{\pi}^{\mathrm{KL}}}(\overline{U_{r}}))=\operatorname{D_{\mathrm{KL}}}(T_{\sharp}\pi||T_{\sharp}\widetilde{\pi}^{\mathrm{KL}}(U_{r}))=\operatorname{D_{\mathrm{KL}}}(\pi||\widetilde{\pi}^{\mathrm{KL}}(U_{r})),$
for any $U_{r}$ such that $U_{r}^{\top}C(\pi)U_{r}=I_{r}$. We conclude that the left-hand side of (D.6) and (2.16) are the same. Furthermore, because $\mathrm{d}\overline{\pi}(\overline{x})=\mathrm{d}\pi(T^{-1}(\overline{x}))|\det\nabla T^{-1}(\overline{x})|$ and $T^{-1}(\overline{x})=C(\pi)^{1/2}\overline{x}+m(\pi)$, we can write
$H(\overline{\pi})=\mathbb{E}_{\overline{\pi}}\left[C(\pi)^{1/2}\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\circ T^{-1}\right)\nabla\ln\left(\frac{\mathrm{d}\pi}{\mathrm{d}x}\circ T^{-1}\right)^{\top}C(\pi)^{1/2}\right]=C(\pi)^{1/2}H(\pi)C(\pi)^{1/2}.$
We conclude that the right-hand side of (D.6) and (2.16) are the same. Finally, evaluating (2.16) at $U_{r}=U_{r}^{\downarrow}$ yields (2.17) and concludes the proof. ∎
D.5 Proof of Theorem 9
Proof.
By direct application of Theorem 4 we have the relation
$\operatorname{D_{\mathrm{KL}}}(\pi^{Y}\,||\,\widetilde{\pi}^{Y}(V_{r}))\leq\frac{1}{2}\operatorname{tr\,}(V_{\perp}^{\top}M^{Y}V_{\perp})-\frac{d-r}{2}+\frac{1}{2}\ln\det\left(V_{\perp}^{\top}(2I_{d}+H^{Y}-M^{Y})V_{\perp}\right),$
where $M^{Y}:=\mathbb{E}_{X\sim\pi^{Y}}[XX^{\top}]$ and $H^{Y}:=\mathbb{E}_{X\sim\pi^{Y}}[\nabla\ln\ell(X)\nabla\ln\ell(X)^{\top}]$ denote matrices which are measurable with respect to the random variable $Y$. It remains to to integrate both sides of the inequality with respect to the law $\rho$ of this random variable. However, recognizing that $\mathbb{E}_{Y\sim\rho}[M^{Y}]=I_{d}$ and $\mathbb{E}_{Y\sim\rho}[H^{Y}]=H^{\mathrm{df}}$, applying Jensen’s inequality to the concave matrix-valued function $S\mapsto\ln\det S$ concludes the proof. ∎
D.6 Proof of Theorem 13
The following lemma is required for the proof of Theorem 13.
Lemma 16.
For all $\mu$-integrable functions $f$ and $g$ we have the inequality $\frac{\mathbb{E}_{X\sim\mu}[f(X)]^{2}}{\mathbb{E}_{X\sim\mu}[g(X)]}\leq\mathbb{E}_{X\sim\mu}\left[\frac{f(X)^{2}}{g(X)}\right].$
Proof.
This follows from Jensen’s inequality as $(x,y)\mapsto\frac{x^{2}}{y}$ is jointly convex [^4] A bivariate function $(x,y)\mapsto f(x,y)$ is jointly convex if for all $\lambda\in[0,1]$, and for all pairs $(x_{1},y_{1})$ and $(x_{2},y_{2})$, it satisfies the inequality $f(\lambda x_{1}+(1-\lambda)x_{2},\lambda y_{1}+(1-\lambda)y_{2})\leq\lambda f(x_{1},y_{1})+(1-\lambda)f(x_{2},y_{2})$. (see [15, §3.2.6]). ∎
The proof of Theorem 13 then proceeds as follows.
Proof.
Without loss of generality, assume $\ell:\mathbb{R}^{d}\to\mathbb{R}^{+}$ is normalized so that $\mathrm{d}\pi=\ell\mathrm{d}\mu$, or equivalently, $\int\ell\mathrm{d}\mu=1$. Given a matrix $U_{r}\in\mathbb{R}^{d\times r}$ with orthonormal columns and $x_{r}\in\mathbb{R}^{r}$, consider the function $f:\mathbb{R}^{d-r}\to\mathbb{R}$ given by
$f(x_{\perp})=\sqrt{\ell(U_{r}x_{r}+U_{\perp}x_{\perp})}$
with $U_{\perp}$ as any orthogonal complement of $U_{r}$. By construction we have $\int f^{2}\mathrm{d}\mu=1$. For this choice of test function, the dimensional Poincaré inequality (3.6) for the $(d-r)$ dimensional Gaussian measure $\mu_{\perp}$ yields
$\int f^{2}\mathrm{d}\mu_{\perp}-\left(\int f\mathrm{d}\mu_{\perp}\right)^{2}\leq\int|\nabla\ln f|^{2}f^{2}\mathrm{d}\mu_{\perp}-\int\frac{(1-\frac{1}{f}\int f\mathrm{d}\mu_{\perp}-\langle\nabla\ln f,x_{\perp}\rangle)^{2}}{d-r+|x_{\perp}|^{2}}f^{2}\mathrm{d}\mu_{\perp}.$ (D.7)
We next integrate the above inequality with respect to $x_{r}$ with law $\mu_{r}$ given by the $r$-dimensional isotropic Gaussian measure. By (3.3), the left-hand side of (D.7) becomes $1-\int\ell_{r}^{\mathrm{Hell}}\mathrm{d}\mu_{r}=1-(1-d_{\mathrm{Hell}}^{2}(\pi\,,\,\tilde{\pi}_{r}^{\mathrm{Hell}}(U_{r})))^{2}$. The first term in right hand of (D.7) becomes $\int|\nabla\ln f|^{2}f^{2}\mathrm{d}\mu=\int\frac{1}{4}|U_{\perp}\nabla\ln\ell|^{2}\mathrm{d}\pi,$ after recognizing that $\mathrm{d}\pi=f^{2}\mathrm{d}\mu$. Meanwhile, applying Lemma 16 to the latter term in the right-hand side of (D.7) shows that
$-\int\frac{(1-\frac{1}{f}\int f\mathrm{d}\mu_{\perp}-\langle\nabla\ln f,x_{\perp}\rangle)^{2}}{d-r+|x_{\perp}|^{2}}\mathrm{d}\pi\leq-\frac{(1-\mathbb{E}_{\pi}[\frac{1}{\sqrt{\ell}}\int\sqrt{\ell}\mathrm{d}\mu_{\perp}]-\frac{1}{2}\mathbb{E}_{\pi}[\langle U_{\perp}^{\top}\nabla\ln\ell,U_{\perp}^{\top}x\rangle)^{2}}{d-r+\mathbb{E}_{\pi}[|U_{\perp}^{\top}x|^{2}]}.$ (D.8)
This can be further simplified, first by noting that
[…]
and secondly by noting that
$\int\nabla\ln\ell(x)x^{\top}\mathrm{d}\pi(x)=\int xx^{\top}\mathrm{d}\pi(x)-I,$
from applying Stein’s identity (i.e., integration by parts) with the standard Gaussian $\mu$. ∎
Acknowledgments
ML and YM thank Brent Wheelock, Dimitar Trenev, Kevin Daly, Grant Seastream, and Tuan Tran at ExxonMobil for the code and discussions related to Bayesian inference for geophysical applications with generative modeling priors. ML also thanks Nicolas Boumal for discussions on the ManOpt forum related to optimization on matrix manifolds.
Funding
ML and YMM acknowledge support from the US Department of Energy, Office of Advanced Scientific Computing Research, under grant DE-SC0023187 and from the ExxonMobil Technology and Engineering Company. YMM further acknowledges support from the US Department of Energy, Office of Advanced Scientific Computing Research, under grants DE-SC0021226 and DE-SC0023188. TC is supported by the ARC Discovery Project DP210103092. OZ acknowledges support from the ANR JCJC project MODENA (ANR-21-CE46-0006-01).